### v1

# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

Abstract
The forecasting of irregular multivariate time series (IMTS) is crucial in key areas such as healthcare, biomechanics, climate science, and astronomy. However, achieving accurate and practical predictions is challenging due to two main factors. First, the inherent irregularity and data missingness in irregular time series make modeling difficult. Second, most existing methods are typically complex and resource-intensive. In this study, we propose a general framework called APN to address these challenges. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner. Additionally, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Experimental results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao et al., 2018; Brouwer et al., 2019; Kidger et al., 2020; Ham et al., 2019; Vio et al., 2013). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and data missingness (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is: how to design an adaptive patching approach that can adapt to the local information density variations of IMTS and capture complete semantic information.
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural ODEs-based models (Chen et al., 2018; Rubanova et al., 2019; Gravina et al., 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin, 2021; Chen et al., 2018). GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex
Figure 1: Fixed Patching vs. Adaptive Patching.
graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024a), and CycleNet (Lin et al., 2024b) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is: how to design an effective model for IMTSF.
To address the above challenges, we propose a general framework called APN. Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner. Based on the representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency. Finally, predictions are made by a shallow MLP. Results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
The contributions of our paper are summarized as follows:
  • To address the IMTSF, we propose a general framework named APN. This framework leverages adaptive patching to generate high-quality and regular initial patch representations. Based on these representations, we employ a simple query module to integrate contextual information, ensuring the effective design of the framework.
  • We design a novel TAPA module to achieve adaptive patching. By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequence into high-quality, regularized representations in a channel-independent manner, effectively adapting to local variations in information density and capturing complete semantics.
  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
2 Related Work
2.1 Irregular Multivariate Time Series Forecasting
IMTSF is crucial for key domains such as healthcare and climate science. The inherent characteristics of IMTS, such as non-uniform sampling intervals and asynchronous channels, present significant challenges to regular time series forecasting models. To address these characteristics, researchers have proposed various IMTSF models. Some models employ approaches based on continuous-time dynamics (Chen et al., 2018; Rubanova et al., 2019; Schirmer et al., 2022; Brouwer et al., 2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points. For instance, Neural Flows (Bilos et al., 2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers. GRU-ODE-Bayes (Brouwer et al., 2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism. Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al., 2024; Zhang et al., 2024; Liu et al., 2025) to capture complex dependencies in IMTS. For example, GraFITi (Yalavarthi et al., 2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs. T-PatchGNN (Zhang et al., 2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
Segmenting time series into “patches” to capture local patterns has demonstrated remarkable success in regular time series forecasting tasks, as evidenced by models like PatchTST (Nie et al., 2023) and Crossformer (Wang et al., 2022). Inspired by this, the patch-based approach has also been applied to IMTSF. T-PatchGNN (Zhang et al., 2024) employs its “Transformable Patch” mechanism to convert irregular sequences into patches with uniform temporal resolution but variable lengths, thereby maintaining temporal alignment while adapting to the sparsity and density variations of observation points. Meanwhile, TimeCHEAT (Liu et al., 2025) divides IMTS into fixed-length patches and utilizes graph learning within each patch for fine-grained temporal embedding. Both methods ingeniously leverage the patch strategy to transform raw IMTS data into structured representations that are more amenable to model processing, laying the groundwork for subsequent sequence modeling.
Unlike the fixed patching methods mentioned above, this study introduces adaptive patching. By learning dynamically adjustable patch boundaries and employing a time-aware weighted averaging strategy, our method transforms the original irregular sequence into a high-quality regularized patch representation in a channel-independent manner. This approach effectively adapts to local variations in information density and captures complete semantic information.
3 Methodology
Figure 2: The overall framework of APN, which initially divides each univariate irregular time series into a series of unfixed patches using the Adaptive Patching Module. Then the Weighted Aggregation Module generates high-quality and regular initial patch representations. Based on the representations, the Query-based Aggregation Module is utilized to incorporate contextual information. Finally, the Forecasting Decoder outputs the final forecasting results. The Adaptive Patching Module and Weighted Aggregation Module collectively form the Time-Aware Patch Aggregation Module.
3.1 Problem Definition
An Irregular Multivariate Time Series (IMTS) $\mathcal{O}$ consists of $N$ independent univariate sequences $\{o^{n}_{1:L_{n}}\}_{n=1}^{N}$. Each sequence $o^{n}_{1:L_{n}}$ comprises $L_{n}$ observations $(t_{i}^{n},v_{i}^{n})$, where the intervals between timestamps $t_{i}^{n}$ are irregular, and sampling across different variables is typically asynchronous. The IMTS forecasting task is defined as: given historical observations $\mathcal{O}$ and a set of prediction queries for all variables $\mathcal{Q}=\{[q_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$, construct and optimize a prediction model $\mathcal{F}(\cdot)$ capable of accurately predicting the future observed values $\hat{\mathcal{V}}=\{[\hat{V}_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$ corresponding to each query $q_{j}^{n}$, i.e.:
$\mathcal{F}(\mathcal{O},\mathcal{Q})\rightarrow\hat{\mathcal{V}}$ (1)
3.2 Framework Overview
The overall pipeline of APN is shown in Figure 2. Its core component is the Time-Aware Patch Aggregation (TAPA) Module–see Section 3.3, which processes the multivariate irregular sequence in a Channel-Independent manner. Compared with conventional fixed segmentation, TAPA utilizes the Adaptive Patching and Weighted Aggregation. This design explicitly aims to adapt to the local variations in information density inherent in IMTS and to capture complete semantic information. Then, a concise and efficient Query-based Aggregation module–see Section 3.4 generates a representation of historical information for each variable. Finally, this representation is fed into a simple MLP-based Forecasting decoder to make predictions–see Section 3.5. In brief, the core advantage of APN stems from the high-quality representations produced by TAPA, which effectively compresses the historical information through adapting to variations in the distribution and concentration of information and preserving complete semantic context, thus APN achieves state-of-the-art performance with a subsequent simple MLP decoder.
3.3 Time-Aware Patch Aggregation (TAPA)
Traditional Fixed-Span Patching (Zhang et al., 2024; Liu et al., 2025), due to the fixed patch boundaries or predefined time spans, struggles to adapt to local variations in the underlying information content and its temporal distribution (referred to as information density), and to capture complete semantic information in IMTS. To overcome this limitation, the TAPA mechanism devises the Adaptive Patching and Weighted Aggregation to (1) adapt the patching strategy to the local information density variations within each time series, and (2) ensure that crucial semantic information is holistically captured within the generated patches by focusing on the information most important for the prediction task. It efficiently integrates the information in each univariate irregular sequence $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$ by transforming it into regularized patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$. This process is performed independently for each variable, accommodating their unique sampling patterns and information distributions.
3.3.1 Adaptive Patching
The Adaptive Patching works through dynamically learning the temporal coverage range $[t_{p}^{left,n},t_{p}^{right,n}]$ for each target patch. This dynamic learning of boundaries is the primary mechanism by which TAPA adapts to local information density variations. For instance, in sparsely observed regions that nonetheless contain critical information, the learned patch width (controlled by $\lambda_{p}^{n}$) might expand to ensure this information is captured, or its position (controlled by $\delta_{p}^{n}$) might shift to center on these information-rich sparse points. Conversely, in densely observed regions with redundant information, a patch might become narrower to focus only on the novel or most salient signals, thus avoiding dilution by less informative data. Furthermore, this adaptive process enables the model to determine the content of each patch based on the inherent importance of the data, which is crucial for capturing complete semantic information. Specifically, TAPA learns the left boundary offset $\delta_{p}^{n}$ and a log-scale parameter $\lambda_{p}^{n}$ representing the patch width:
$t_{p}^{left,n}$ $=c_{p}-\frac{S_{init}}{2}+\delta_{p}^{n},$ (2)
$t_{p}^{right,n}$ $=t_{p}^{left,n}+\exp(\lambda_{p}^{n}),$ (3)
where $\delta_{p}^{n}$ and $\lambda_{p}^{n}$ are learnable parameters, $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ denote the reference center points and initial average patch width. For observations highly important for prediction (i.e., regions with high information density relevant to the task), different patch boundaries are adjusted to accurately cover these crucial regions, ensuring their semantic integrity. While for regions with noise or low information relevance (low information density), boundaries may contract or shift to reduce their coverage, thus preserving the clarity of the extracted semantic information. The dynamic adjustment of boundaries thus constitutes an effective mechanism for adaptively selecting useful information and forming semantically coherent patches based on learned information importance.
3.3.2 Weighted Aggregation
To fully utilize temporal information from observations, the Learnable Time Embedding $TE(t_{i})\in\mathbb{R}^{D_{te}}$ is first generated for $i$-th timestamp, which consists of a linear layer for scale and a sine-activated linear layer for periodicity. Then the Learnable Time Embedding is concatenated with the original observed value $v_{i}$ to form an augmented representation $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
After determining the dynamic coverage range $[t_{p}^{left,n},t_{p}^{right,n}]$ (which has already adapted to local information densities and aimed to cover semantic regions), we employ a query-like Weighted Aggregation. This weighting mechanism synergizes with the adaptive boundaries to further refine the captured semantic information by emphasizing the most salient points within each adaptively formed patch. For any observation $i$ (timestamp $t_{i}$) of variable $n$, its relationship $\alpha_{i,p}^{n}$ to target patch $p$ is defined by a soft window function:
$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right),$ (4)
where $\sigma(\cdot)$ is the Sigmoid function, and $\text{Softplus}(\kappa^{n})$ is a temperature scaling factor controlled by a learnable scalar parameter $\kappa^{n}$, regulating the smoothness of the soft window boundary transition. Through learning the $\kappa^{n}$, the observations within the dynamically determined patch boundaries are further weighted, where important observations contributing to high local information density receive larger $\alpha_{i,p}^{n}$ in aggregation, while less relevant observations are down-weighted or filtered out, thus ensuring the completeness and purity of the captured semantic information.
After calculating the relationships between observations and target patches, the aggregated representation $\bar{h}_{p}^{n}$ for the $p$-th patch of variable $n$ can be obtained through a weighted average of the augmented features $\tilde{v}^{n}_{i}$:
$\bar{h}_{p}^{n}=\frac{\sum_{i}\alpha_{i,p}^{n}\cdot\tilde{v}^{n}_{i}}{\sum_{i}\alpha_{i,p}^{n}+\epsilon}\in\mathbb{R}^{1+D_{te}},$ (5)
where $\epsilon$ is a small positive constant to prevent division by zero. Subsequently, we transform $\bar{h}_{p}^{n}\in\mathbb{R}^{(1+D_{te})}$ into the uniform hidden representations $h_{p}^{n}\in\mathbb{R}^{D}$ to enhance the expressiveness through a linear projection layer $h_{p}^{n}=\text{Linear}_{D}(\bar{h}_{p}^{n})$.
Through the aforementioned mechanisms, each univariate irregular sequence $X^{n}$ is efficiently transformed into a refined and structurally regularized patch representation sequence $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$.
3.4 Query-based Weighted Aggregation
After TAPA generates regularized patch representation sequences $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$ for each variable, APN employs a concise mechanism to summarize them into a single representation $H_{c}^{n}$, obviating complex inter-sequence transformation modules. To preserve the sequential order of patches, standard positional encodings $PE\in\mathbb{R}^{P\times D}$ are added to $H^{n}$, yielding position-aware patch representations $H_{pe}^{n}=H^{n}+PE$. Then, to dynamically select and fuse information most relevant to the prediction task, a learnable query vector $q^{n}\in\mathbb{R}^{D}$ is introduced for each variable $n$. This query vector interacts with all $P$ position-aware patches $H_{pe}^{n}$, calculating unregularized importance scores $s_{p}^{n}$ via dot product, which are then regularized into weight coefficients $\beta^{n}=[\beta_{1}^{n},\dots,\beta_{P}^{n}]$ using the Softmax function. Finally, the representation $H_{c}^{n}$ of variable $n$ is obtained by a weighted sum of all patches $h_{pe,p}^{n}$:
$s_{p}^{n}$ $=\frac{q^{n}\cdot(h_{pe,p}^{n})^{T}}{\sqrt{D}}\quad\text{for }p=1,\dots,P,$ (6)
$\beta^{n}$ $=\text{Softmax}(\{s_{p}^{n}\}_{p=1}^{P}),$ (7)
$H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n},$ (8)
where the summary representations of all variables $[H_{c}^{1},\dots,H_{c}^{N}]$ form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$, which is then processed by a layer normalization to stabilize subsequent MLP input. This learnable Query-based Weighted Aggregation is concise and efficient, avoiding complex key-value projections or pairwise inter-patch interactions, demonstrating APN’s simple yet effective design.
3.5 Forecasting Decoder
After obtaining the summary historical representation $H_{final}^{n}\in\mathbb{R}^{D}$ for each variable $n$ and the target query time point $\tau_{k}$, APN employs a two-layer MLP decoder for prediction. The decoder takes the concatenation of the variable’s summary representation $H_{final}^{n}$ and the learnable temporal encoding $TE(\tau_{k})$ of the query time point $\tau_{k}$ as input, outputting the predicted value $\hat{v}^{n}_{k}$:
$\hat{v}^{n}_{k}=\text{MLP}(\text{Concat}(H_{final}^{n},TE(\tau_{k})))\in\mathbb{R}$ (9)
3.6 Loss Function
The overall training objective of the APN model is to minimize the Mean Squared Error (MSE) between predicted and true observed values. For given historical observations $\mathcal{O}_{hist}$ and a set of query time points $\mathcal{T}_{pred}=\{\tau_{k}\}_{k=1}^{H}$, the loss function $\mathcal{L}$ is defined as:
$\mathcal{L}=\frac{1}{N\cdot H}\sum_{n=1}^{N}\sum_{k=1}^{H}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
where $v^{n}_{k}$ is the true value of variable $n$ at query time $\tau_{k}$, and $\hat{v}^{n}_{k}$ is the model’s prediction.
4 Experiments
4.1 Setup
Datasets: To comprehensively evaluate the model’s performance, we select four widely-used IMTS datasets, including PhysioNet, MIMIC, Activity, and USHCN. These datasets span multiple domains such as healthcare, biomechanics, and climate science. More details of the benchmark datasets are included in Appendix A.1.
Baselines: To comprehensively evaluate the performance of APN, we compare it with nineteen baseline models. These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), Crossformer (Wang et al., 2022), Graph Wavenet (Wu et al., 2019), MTGNN (Wu et al., 2020), StemGNN (Cao et al., 2020), CrossGNN (Huang et al., 2023), and FourierGNN (Yi et al., 2023). (2) IMTS classification/imputation models: including GRU-D (Che et al., 2018), SeFT (Horn et al., 2020), RainDrop (Zhang et al., 2022), Warpformer (Zhang et al., 2023), and mTAND (Shukla and Marlin, 2021). (3) IMTSF models: including Latent ODEs (Rubanova et al., 2019), CRU (Schirmer et al., 2022), and Neural Flows (Bilos et al., 2021), T-PATCHGNN (Zhang et al., 2024), and TimeCHEAT (Liu et al., 2025). For detailed information on the baseline models, please refer to Appendix A.2.
Implementation Details: We implement APN and all related experiments using the PyTorch framework on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (@ 2.90GHz) and an NVIDIA GeForce RTX 3090 GPU. For our APN model, we set the hidden dimension $D$ according to different datasets. Specifically, the hidden dimension $D$ is set to 64 for the PhysioNet and USHCN, while it is set to 16 for the Human Activity and MIMIC. The time encoding dimension $D_{te}$ is uniformly set to 10. The number of patches $P$ is adjusted based on the characteristics of the datasets: 20 for PhysioNet, 16 for MIMIC, 75 for Human Activity, and 4 for USHCN. The training batch size is set to 256 for all datasets. The model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-2}$. We train the model for a maximum of 200 epochs and employ an early stopping strategy with a patience of 50. All experiments are repeated 5 times, and the mean and standard deviation of the performance metrics are reported to ensure the robustness of the results.
Table 1: Overall performance evaluated by MSE and MAE (mean $\pm$ std). The best-performing and second-best results are highlighted in bold and underline, respectively.
Dataset PhysioNet MIMIC Human Activity USHCN
Metric MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-2}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-1}$ MAE×$\times 10^{-1}$
DLinear 41.86 $\pm$ 0.05 15.52 $\pm$ 0.03 4.90 $\pm$ 0.00 16.29 $\pm$ 0.05 4.03 $\pm$ 0.01 4.21 $\pm$ 0.01 6.21 $\pm$ 0.00 3.88 $\pm$ 0.02
TimesNet 16.48 $\pm$ 0.11 6.14 $\pm$ 0.03 5.88 $\pm$ 0.08 13.62 $\pm$ 0.07 3.12 $\pm$ 0.01 3.56 $\pm$ 0.02 5.58 $\pm$ 0.05 3.60 $\pm$ 0.04
PatchTST 12.00 $\pm$ 0.23 6.02 $\pm$ 0.14 3.78 $\pm$ 0.03 12.43 $\pm$ 0.10 4.29 $\pm$ 0.14 4.80 $\pm$ 0.09 5.75 $\pm$ 0.01 3.57 $\pm$ 0.02
Crossformer 6.66 $\pm$ 0.11 4.81 $\pm$ 0.11 2.65 $\pm$ 0.10 9.56 $\pm$ 0.29 4.29 $\pm$ 0.20 4.89 $\pm$ 0.17 5.25 $\pm$ 0.04 3.27 $\pm$ 0.09
Graph Wavenet 6.04 $\pm$ 0.28 4.41 $\pm$ 0.11 2.93 $\pm$ 0.09 10.50 $\pm$ 0.15 2.89 $\pm$ 0.03 3.40 $\pm$ 0.05 5.29 $\pm$ 0.04 3.16 $\pm$ 0.09
MTGNN 6.26 $\pm$ 0.18 4.46 $\pm$ 0.07 2.71 $\pm$ 0.23 9.55 $\pm$ 0.65 3.03 $\pm$ 0.03 3.53 $\pm$ 0.03 5.39 $\pm$ 0.05 3.34 $\pm$ 0.02
StemGNN 6.86 $\pm$ 0.28 4.76 $\pm$ 0.19 1.73 $\pm$ 0.02 7.71 $\pm$ 0.11 8.81 $\pm$ 0.37 6.90 $\pm$ 0.02 5.75 $\pm$ 0.09 3.40 $\pm$ 0.09
CrossGNN 7.22 $\pm$ 0.36 4.96 $\pm$ 0.12 2.95 $\pm$ 0.16 10.82 $\pm$ 0.21 3.03 $\pm$ 0.10 3.48 $\pm$ 0.08 5.66 $\pm$ 0.04 3.53 $\pm$ 0.05
FourierGNN 6.84 $\pm$ 0.35 4.65 $\pm$ 0.12 2.55 $\pm$ 0.03 10.22 $\pm$ 0.08 2.99 $\pm$ 0.02 3.42 $\pm$ 0.02 5.82 $\pm$ 0.06 3.62 $\pm$ 0.07
GRU-D 5.59 $\pm$ 0.09 4.08 $\pm$ 0.05 1.76 $\pm$ 0.03 7.53 $\pm$ 0.09 2.94 $\pm$ 0.05 3.53 $\pm$ 0.06 5.54 $\pm$ 0.38 3.40 $\pm$ 0.28
SeFT 9.22 $\pm$ 0.18 5.40 $\pm$ 0.08 1.87 $\pm$ 0.01 7.84 $\pm$ 0.08 12.20 $\pm$ 0.17 8.43 $\pm$ 0.07 5.80 $\pm$ 0.19 3.70 $\pm$ 0.11
RainDrop 9.82 $\pm$ 0.08 5.57 $\pm$ 0.06 1.99 $\pm$ 0.03 8.27 $\pm$ 0.07 14.92 $\pm$ 0.14 9.45 $\pm$ 0.05 5.78 $\pm$ 0.22 3.67 $\pm$ 0.17
Warpformer 5.94 $\pm$ 0.35 4.21 $\pm$ 0.12 1.73 $\pm$ 0.04 7.58 $\pm$ 0.13 2.79 $\pm$ 0.04 3.39 $\pm$ 0.03 5.25 $\pm$ 0.05 3.23 $\pm$ 0.05
mTAND 6.23 $\pm$ 0.24 4.51 $\pm$ 0.17 1.85 $\pm$ 0.06 7.73 $\pm$ 0.13 3.22 $\pm$ 0.07 3.81 $\pm$ 0.07 5.33 $\pm$ 0.05 3.26 $\pm$ 0.10
Latent-ODE 6.05 $\pm$ 0.57 4.23 $\pm$ 0.26 1.89 $\pm$ 0.19 8.11 $\pm$ 0.52 3.34 $\pm$ 0.11 3.94 $\pm$ 0.12 5.62 $\pm$ 0.03 3.60 $\pm$ 0.12
CRU 8.56 $\pm$ 0.26 5.16 $\pm$ 0.09 1.97 $\pm$ 0.02 7.93 $\pm$ 0.19 6.97 $\pm$ 0.78 6.30 $\pm$ 0.47 6.09 $\pm$ 0.17 3.54 $\pm$ 0.18
Neural Flow 7.20 $\pm$ 0.07 4.67 $\pm$ 0.04 1.87 $\pm$ 0.05 8.03 $\pm$ 0.19 4.05 $\pm$ 0.13 4.46 $\pm$ 0.09 5.35 $\pm$ 0.05 3.25 $\pm$ 0.05
T-PATCHGNN 4.98 $\pm$ 0.08 3.72 $\pm$ 0.03 1.69 $\pm$ 0.03 7.22 $\pm$ 0.09 2.66 $\pm$ 0.03 3.15 $\pm$ 0.02 5.00 $\pm$ 0.04 3.08 $\pm$ 0.04
TimeCHEAT 5.03 $\pm$ 0.07 3.88 $\pm$ 0.03 1.71 $\pm$ 0.02 7.41 $\pm$ 0.10 4.05 $\pm$ 0.10 4.64 $\pm$ 0.07 4.25 $\pm$ 0.07 3.18 $\pm$ 0.05
APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.77 $\pm$ 0.00 3.10 $\pm$ 0.01 4.09 $\pm$ 0.06 2.92 $\pm$ 0.05
4.2 Main results
We compare the APN model with nineteen baseline models on four challenging datasets—see Table 1. We have the following observations: 1) APN achieves leading prediction accuracy across the board. On all datasets, APN achieves optimal or highly competitive forecasting performance. Compared to the second-best performing model, T-PATCHGNN, APN achieves significant reductions in MSE and MAE metrics by approximately 9.63% and 5.24%, respectively. 2) APN demonstrates exceptional cross-domain generalization capability and robustness. Whether on IMTS datasets with different characteristics in healthcare (PhysioNet, MIMIC), biomechanics (Human Activity), or climate science (USHCN), APN consistently performs excellently. The outstanding performance of APN can be attributed to its innovative use of adaptive patching, which generates high-quality and regular initial patch representations. Building on these representations, APN employs a streamlined query module to effectively integrate contextual information, ensuring a lightweight and efficient framework design.
4.3 Ablation Studies
To validate the contributions of each key component of our APN model, we conduct ablation experiments—see Table 2. We have the following observations: 1) w/o Adaptive Patching: Replacing the learnable boundaries in TAPA with fixed time windows leads to performance degradation across all datasets, demonstrating the importance of the adaptive boundary learning mechanism. 2) w/o Weighted Aggregation: Replacing the time membership-based weighted averaging in TAPA with a simple arithmetic average of observations within the boundaries results in a significant performance drop. This highlights the core value of the time membership weighting strategy in synergizing with dynamic boundaries to further refine information and balance the information density within patches. 3) w/o Query-Based Patch Aggregation: Replacing the query-based intra-variable patch aggregation module with a simple linear mapping also leads to a notable performance loss, indicating that even with high-quality initial patch representations, effective context information integration remains crucial.
Table 2: Ablation studies for APN in terms of the best result highlighted in bold.
Dataset PhysioNet MIMIC Human Activity USHCN
Metric MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-2}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-1}$ MAE×$\times 10^{-1}$
w/o Adaptive Patching 4.58 $\pm$ 0.05 3.45 $\pm$ 0.03 1.65 $\pm$ 0.02 6.94 $\pm$ 0.04 2.78 $\pm$ 0.01 3.11 $\pm$ 0.00 4.38 $\pm$ 0.05 3.07 $\pm$ 0.04
w/o Weighted Aggregation 9.92 $\pm$ 0.11 5.81 $\pm$ 0.07 1.64 $\pm$ 0.03 6.96 $\pm$ 0.05 4.36 $\pm$ 0.12 4.28 $\pm$ 0.03 4.19 $\pm$ 0.02 2.95 $\pm$ 0.03
w/o Query-based Patch Aggregation 9.92 $\pm$ 0.10 4.71 $\pm$ 0.09 3.57 $\pm$ 0.03 11.58 $\pm$ 0.10 2.81 $\pm$ 0.02 3.18 $\pm$ 0.05 4.51 $\pm$ 0.04 3.52 $\pm$ 0.02
APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.77 $\pm$ 0.00 3.10 $\pm$ 0.01 4.09 $\pm$ 0.06 2.92 $\pm$ 0.05
4.4 Parameter Sensitivity
Figure 3: Parameter sensitivity studies of main hyper-parameters in APN
To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Table 3. We have the following observations: 1) From Figures 3(a) and 3(b), we observe that for most datasets, a larger $P$ does not necessarily lead to better performance. This suggests that an appropriate $P$ value can better balance the capture of local information and model complexity. 2) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$). In general, a smaller hidden dimension may already be sufficient to capture effective information, while an excessively large dimension can sometimes lead to a slight performance decline. This indicates that the choice of $D$ should be comprehensively considered based on the scale and sparsity of the dataset. 3) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$). The results indicate that a moderate $D_{te}$ value can provide the model with effective temporal information. An excessively high time encoding dimension may introduce too much noise, interfering with the learning of core features. 4) The APN model exhibits a certain sensitivity to the selection of these core hyperparameters, but it generally achieves competitive results within a reasonable range. In practical applications, appropriate tuning should be performed based on the specific characteristics of the dataset.
4.5 Scalability and Efficiency Analysis
Figure 4: We evaluate the memory usage (GB), number of parameters (M), average training time per epoch (s), and total inference time (s) of APN and three representative baselines for IMTSF. All experiments are conducted on the PhysioNet dataset with a consistent batch size of 32 to ensure fair comparison. A lower value indicates better performance.
To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu et al., 2025), CRU (Schirmer et al., 2022), T-PatchGNN (Zhang et al., 2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4. The results demonstrate that APN exhibits significant advantages across all key efficiency metrics. This is attributed to its core TAPA module, which efficiently generates information-condensed and regularized initial representations. These representations support a highly streamlined information aggregation and prediction architecture, ultimately achieving a lightweight and highly efficient model design.
4.6 Varying Horizon
Table 3: Performance of varying observation and forecast horizons.
Algorithm History=3h, Forecast=45h History=12h, Forecast=36h History=36h, Forecast=12h History=45h, Forecast=3h
MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$
DLinear 51.82 17.13 43.56 15.73 41.63 15.48 41.23 15.51
TimesNet 57.30 10.70 24.95 7.62 13.57 5.50 13.86 5.65
PatchTST 42.18 13.67 18.56 7.80 9.85 5.11 8.53 4.64
Crossformer 9.48 5.86 8.57 5.70 5.70 4.47 5.33 4.44
Graph Wavenet 9.43 5.86 7.23 4.82 4.71 3.90 4.10 3.73
MTGNN 9.83 5.95 7.48 5.01 5.08 3.99 5.22 4.19
StemGNN 8.70 5.37 7.46 4.84 6.65 4.69 5.47 4.56
CrossGNN 10.44 6.56 7.97 5.37 6.87 4.73 4.80 4.24
FourierGNN 9.59 5.61 7.95 4.99 6.35 4.61 5.37 4.34
GRU-D 8.18 4.99 6.89 4.55 4.42 3.66 4.44 3.79
SeFT 9.78 5.55 9.30 5.41 9.15 5.15 8.76 5.57
RainDrop 10.47 5.72 9.89 5.62 9.70 5.40 9.28 5.62
Warpformer 8.48 5.13 7.57 4.83 5.60 4.09 6.44 4.67
mTAND 8.45 5.23 7.11 4.67 5.71 4.17 5.44 4.33
Latent-ODE 8.25 5.04 7.20 4.69 6.70 4.36 7.10 5.33
CRU 9.20 5.38 9.20 5.31 9.50 5.41 11.60 6.98
Neural Flow 8.30 4.99 8.50 5.27 7.70 4.68 7.40 5.10
T-PATCHGNN 8.01 4.87 6.48 4.19 4.14 3.31 3.69 3.25
TimeCHEAT 9.43 5.41 6.64 4.57 5.26 4.04 2.79 3.35
APN (ours) 8.79 4.90 5.97 3.94 3.91 3.08 2.51 2.65
To further evaluate the robustness of APN under varying combinations of observation history lengths and forecasting horizons, we conduct experiments on the PhysioNet dataset—as shown in Table 3. The results demonstrate that APN exhibits strong predictive capabilities and broad adaptability across multiple "History-Forecast" settings. In scenarios where historical information is relatively sufficient or the prediction span is moderate, APN achieves optimal performance in both MSE and MAE metrics, leading comprehensively. Under extreme settings with extremely limited historical information and long prediction spans, APN’s performance is comparable to several strong baseline models, ranking second only to T-PATCHGNN. Considering the significant challenges faced by all models under such extreme sparsity conditions, APN’s TAPA mechanism, which relies on learning adaptive features from data, may not fully realize its potential when historical information is scarce. Nevertheless, APN still demonstrates competitive robustness.
5 Conclusion
This paper addresses the challenges of IMTS forecasting by proposing a general and efficient APN. The core of APN lies in its novel TAPA module. Through learned dynamically adjustable patch boundaries and a time-aware weighting strategy, TAPA adaptively transforms irregular sequences into a fixed number of high-quality, regularized patch representations with highly condensed information, operating in a channel-independent manner. Leveraging the superior initial representations generated by TAPA, APN employs an exceptionally lean query module to integrate historical information and completes predictions via a shallow MLP. Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.
Appendix A Detailed Experimental Setup
A.1 Dataset Details
To comprehensively and rigorously evaluate the predictive performance of our proposed APN model, we conduct extensive experiments on four real-world Irregular Multivariate Time Series (IMTS) datasets. These datasets are selected for their widespread use in benchmarking and public availability, representing typical characteristics of IMTS across different domains. Table 4 summarizes the key statistical features of these datasets.
For all datasets, samples are randomly partitioned into training, validation, and test sets, adhering to the standard 60%:20%:20% ratio. Regarding data preprocessing schemes (e.g., missing value handling, normalization methods, timestamp conversion) and prediction task configurations (e.g., historical observation window length, future prediction horizon length), we strictly follow established mainstream experimental protocols, particularly referencing benchmark settings established in related works such as T-PatchGNN [Zhang et al., 2024], to ensure the fairness and comparability of our experimental results. Next, we detail the characteristics of each dataset and its specific application in this study.
Table 4: Dataset statistics.
Dataset # Variable Avg # Obs. Avg Length Max Length
PhysioNet 41 10.7 75 216
MIMIC 96 1.5 46.1 643
Human Activity 12 30.2 120.6 130
USHCN 5 36.1 163.9 214
  • PhysioNet [Ansari et al., 2023]: Physionet [^1] https://archive.physionet.org/challenge/2012/ dataset contains time series data collected from 12,000 ICU patients. Each record details 41 clinical physiological indicators during the first 48 hours after ICU admission, with irregular sampling intervals and a significant amount of missing values. Following the standard settings of [Ansari et al., 2023, Rubanova et al., 2019], we use the observation data from the first 24 hours after each patient’s admission to predict their physiological indicator values for the subsequent 24 hours. Original data and access methods are available on the official PhysioNet website.
  • MIMIC [Naiman et al., 2024]: MIMIC [^2] https://physionet.org/content/mimiciii/1.4/ dataset is a large, publicly available database of de-identified critical care medical information. Following the commonly used preprocessing procedures in studies such as [Bilos et al., 2021, Naiman et al., 2024], we filter and process records to obtain 23,457 patient entries. Each record contains 96 irregularly sampled clinical variables from the first 48 hours after patient admission, also exhibiting significant irregularity and missingness. The experimental task is set to utilize data from the first 24 hours to predict variable values within the subsequent 24 hours. Access to this dataset requires certification and adherence to data use agreements via PhysioNet.
  • Human Activity [Zhang et al., 2023]: Human Activity [^3] https://archive.ics.uci.edu/dataset/196/localization+data+for+person+activity dataset comprises 12 motion-related variables (e.g., 3D acceleration, 3D gyroscope) recorded by sensors worn on different body parts (wrist, chest, ankle) of test subjects, with irregular sampling times. To construct IMTS samples suitable for forecasting tasks, we segment the original long sequences into clips of 4000 milliseconds and use the observation data from the first 3000 milliseconds to predict sensor readings for the next 1000 milliseconds. Through this process, we obtain a total of 5,400 IMTS samples to evaluate the model’s capability in biomechanical signal prediction.
  • USHCN [Ansari et al., 2023]: The USHCN dataset [^4] https://www.osti.gov/biblio/1394920 contains daily data for 5 core climate variables (daily maximum temperature, daily minimum temperature, daily precipitation, daily snow depth, daily snowfall) recorded from 1,114 weather stations across the United States between 1996 and 2000. This dataset is characterized by its long time span and naturally occurring spatial and temporal irregularities and data missingness due to station maintenance, equipment failure, etc. Following the experimental setup of [Ansari et al., 2023, Rubanova et al., 2019], we process the data for the task of using climate records from the past 24 months to predict climate conditions for the next 1 month, generating a total of 26,736 IMTS samples to test the model’s predictive performance on long-period, sparse climate data.
A.2 Baseline Model Details
To construct a comprehensive and convincing benchmark for the IMTS forecasting task, we integrate nineteen relevant baseline models for meticulous comparison. In selecting and evaluating these baseline models, we adhere to established mainstream experimental protocols [Zhang et al., 2024]. Specifically, we adapt state-of-the-art models from regular multivariate time series (RMTS) forecasting, IMTS classification, and IMTS imputation for the current IMTS forecasting task by appropriate modifications (e.g., altering the output layer for prediction tasks and incorporating timestamps and mask information as input features).
To ensure fairness and reliability of comparison, for some baseline models that have been rigorously evaluated in recent related works such as T-PatchGNN [Zhang et al., 2024], we prioritize referencing and reporting their publicly available optimal experimental results. For other models, or when evaluation under specific experimental settings is required, we reproduce them based on their official open-source implementations within a unified and consistent experimental framework. During reproduction, we conduct exploratory searches around the hyperparameter configurations recommended in the original papers and select the optimal parameter combinations based on performance on the validation set to report their best performance. This approach, combining references to existing SOTA results with careful reproduction, aims to provide a broad yet rigorous performance reference system for our APN model.
Next, we detail the core ideas, key mechanisms, and reference code sources for each baseline model:
Regular MTSF models:
  • DLinear [Zeng et al., 2023]: Models the trend and periodicity/residual terms of a time series using two independent single-layer linear networks, respectively, and then sums them to obtain the final prediction. Its design philosophy is that simple and effective linear models remain competitive in many forecasting tasks. Code Reference: https://github.com/cure-lab/LTSF-Linear
  • TimesNet [Wu et al., 2023]: Innovatively transforms one-dimensional time series into two-dimensional tensors to capture multi-periodicity. It identifies major periods via Fourier transform and reshapes sub-sequences of different periods into 2D tensors (with period-based "columns" and intra-period "rows"), then uses a 2D convolutional network to extract features, aiming to uniformly handle series with complex periodicities. Code Reference: https://github.com/thuml/TimesNet
  • PatchTST [Nie et al., 2023]: Segments each univariate time series into fixed-length non-overlapping patches, which serve as input tokens for a Transformer. It employs a Channel Independence strategy, modeling each variable independently, and learns temporal dependencies through a shared Transformer backbone. Code Reference: https://github.com/yuqinie/PatchTST
  • Crossformer [Wang et al., 2022]: A Transformer variant with a two-stage attention mechanism. The first stage captures intra-sequence temporal dependencies through dimension embedding and temporal self-attention; the second stage explicitly models interactions between different variables (dimensions) via an innovative Dimension-Segment-Wise Attention. Code Reference: https://github.com/Thinklab-AI/Crossformer
  • Graph Wavenet [Wu et al., 2019]: Combines Graph Convolutional Networks (GCN) with stacked dilated causal convolutions. It uses an adaptive adjacency matrix to learn dynamic spatial dependencies between variables and efficiently captures long-range temporal dependencies through dilated convolutions. Code Reference: https://github.com/nnzhan/Graph-WaveNet
  • MTGNN [Wu et al., 2020]: Proposes a general framework comprising a graph learning module, a graph convolution module, and a temporal convolution module. Its graph learning module adaptively infers unidirectional relationships between variables through node embedding and metric learning. Code Reference: https://github.com/nnzhan/MTGNN
  • StemGNN [Cao et al., 2020]: First transforms spatio-temporal graph data into the spectral domain (via Graph Fourier Transform and Discrete Fourier Transform), then designs modules in the spectral domain to capture spatial and temporal dependencies, and finally obtains predictions through inverse transformation. Code Reference: https://github.com/microsoft/StemGNN
  • CrossGNN [Huang et al., 2023]: To handle noise in multivariate time series, this model first generates multiple time series copies at different noise levels (data augmentation), then performs information interaction and feature refinement via cross-scale GNN and cross-variable GNN. Code Reference: https://github.com/hqh0728/CrossGNN
  • FourierGNN [Yi et al., 2023]: Treats multivariate time series from a pure graph perspective, where each time series is a node on the graph, and different time points are node features. Graph neural operations are performed in the Fourier domain to achieve prediction. Code Reference: https://github.com/aikunyi/FourierGNN
IMTS classification/imputation models:
  • GRU-D [Che et al., 2018]: An improved GRU model that models the influence of observation intervals through an introduced learnable time decay term and handles missing values via feature-level imputation to dynamically update hidden states. Code Reference: https://github.com/zhiyongc/GRU-D
  • SeFT (Set Functions for Time Series) [Horn et al., 2020]: Treats observation points of an irregular time series (typically as (time, value, variable ID) tuples) as an unordered set and models them using set functions like Deep Sets, naturally handling irregularity and variable length. Code Reference: https://github.com/mims-harvard/Raindrop
  • RainDrop [Zhang et al., 2022]: Models irregularly sampled data from a sensor network as a graph, where each observation point is a node. It captures spatio-temporal dependencies via graph attention networks and temporal attention mechanisms, generating interpolations or predictions for arbitrary query time points. Code Reference: https://github.com/mims-harvard/Raindrop
  • Warpformer [Zhang et al., 2023]: A multi-scale modeling approach for irregular clinical time series. Its core is the introduction of a learnable dynamic time warping module that adaptively aligns time series of different scales and sampling rates to a reference timeline before feeding them into a Transformer for representation learning. Code Reference: https://github.com/imJiawen/Warpformer
  • mTAND [Shukla and Marlin, 2021]: A multi-time attention network designed for irregularly sampled time series. It learns embedding representations for continuous timestamps and generates fixed-length context representations for variable-length input sequences, used for interpolation or classification. Code Reference: https://github.com/reml-lab/mTAN
IMTSF models:
  • Latent ODEs [Rubanova et al., 2019]: Compresses historical observation sequences (potentially variable-length and irregular) into a latent initial state via an RNN encoder, then evolves this state in latent space using a Neural ODE (an ordinary differential equation parameterized by a neural network), and finally generates predictions through a decoder. Code Reference: https://github.com/YuliaRubanova/latent-ode
  • CRU (Continuous Recurrent Units) [Schirmer et al., 2022]: Proposes a continuous-time recurrent unit whose hidden state evolution is driven by a linear stochastic differential equation, with irregularly spaced observation updates handled via Kalman filtering theory. Code Reference: https://github.com/boschresearch/Continuous-Recurrent-Units
  • Neural Flows [Bilos et al., 2021]: Reformulates the solving process of Neural ODEs as a continuous normalizing flow, aiming to improve numerical solution efficiency and stability by directly modeling the solution function of the ODE rather than its derivative. Code Reference: https://github.com/mbilos/neural-flows-experiments
  • T-PATCHGNN [Zhang et al., 2024]: Specifically designed for IMTS forecasting. Its core is to transform irregular univariate sequences into time-aligned patch sequences via a transformable patching method, then combine Transformers to capture intra-patch temporal dependencies and utilize time-adaptive graph neural networks to model dynamic inter-variable dependencies. Code Reference: https://github.com/usail-hkust/t-PatchGNN
  • TimeCHEAT [Liu et al., 2025]: Proposes a channel harmony strategy, employing channel-dependent graph convolutions within patches and channel-independent Transformers between patches to balance information processing across different channels. Code Reference: https://github.com/Alrash/TimeCHEAT
For all baseline models, we endeavor to use the official implementations provided by their authors or widely recognized community implementations. Regarding hyperparameter settings, we first refer to the recommendations in the original papers and then adjust them based on performance on the validation set to ensure fair comparison.
Appendix B Limitation
Although the APN method demonstrates its effectiveness, several aspects warrant further exploration:
  • Limited Scope of Validation Datasets: APN’s current validation primarily focuses on specific types of real-world datasets. Its adaptability and performance on more diverse domains, such as high-frequency financial data, industrial data with strong periodicities, or traffic flow with complex covariates, require further investigation to fully assess its generalization capabilities.
  • Performance in Extreme Scenarios: In situations with exceptionally short observation histories and very long prediction horizons, APN’s performance can be suboptimal. This is potentially because the data-driven adaptive mechanism struggles to converge effectively with extremely scarce information. Future work could explore incorporating prior knowledge or regularization to enhance robustness under such data-sparse conditions.

### v2

@@
 Abstract
-The forecasting of irregular multivariate time series (IMTS) is crucial in key areas such as healthcare, biomechanics, climate science, and astronomy.
-However, achieving accurate and practical predictions is challenging due to two main factors.
-First, the inherent irregularity and data missingness in irregular time series make modeling difficult.
-Second, most existing methods are typically complex and resource-intensive.
-In this study, we propose a general framework called APN to address these challenges.
-Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching.
-By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner.
-Additionally, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency.
-Finally, predictions are made by a shallow MLP.
-Experimental results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
+The forecasting of irregular multivariate time series (IMTS) is a critical task in domains like healthcare and climate science.
+However, this task faces two significant hurdles: 1) the inherent non-uniformity and missing data in IMTS complicate the modeling of temporal dynamics, and 2) existing methods often rely on computationally expensive architectures.
+To address these dual challenges, we introduce APN, a general and efficient forecasting framework.
+At the core of APN is a novel Time-Aware Patch Aggregation (TAPA) module that introduces an aggregation-based paradigm for adaptive patching, moving beyond the limitations of fixed-span segmentation and interpolation-based methods.
+TAPA first learns dynamic temporal boundaries to define data-driven segments.
+Crucially, instead of resampling or interpolating, it directly computes patch representations via a time-aware weighted aggregation of all raw observations, where weights are determined by each observation’s temporal relevance to the segment.
+This approach provides two key advantages: it preserves data fidelity by avoiding the introduction of artificial data points and ensures complete information coverage by design.The resulting regularized and information-rich patch representations enable the use of a lightweight query module for historical context aggregation and a simple MLP for final prediction.
+Extensive experiments on multiple real-world datasets demonstrate that APN establishes a new state-of-the-art, significantly outperforming existing methods in both prediction accuracy and computational efficiency.
 1 Introduction
@@
 Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making.
-However, the inherent irregularity of observations and data missingness (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
+However, the inherent irregularity of observations and missing data (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
 To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), have adopted the Fixed Patching approach—see Figure 1a.
@@
 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context.
-Therefore, the first challenge is: how to design an adaptive patching approach that can adapt to the local information density variations of IMTS and capture complete semantic information.
+Therefore, the first challenge is to design an adaptive patching approach that can effectively adapt to local information density variations in IMTS and capture the complete semantic context.
+Figure 1: Fixed Patching vs. Adaptive Patching.
 Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times.
 For example, neural ODEs-based models (Chen et al., 2018; Rubanova et al., 2019; Gravina et al., 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin, 2021; Chen et al., 2018).
-GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex
-Figure 1: Fixed Patching vs. Adaptive Patching.
-graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024).
+GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024).
 These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS.
 While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios.
-Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024a), and CycleNet (Lin et al., 2024b) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters.
-Therefore, the second challenge is: how to design an effective model for IMTSF.
+Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024b), and CycleNet (Lin et al., 2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters.
+Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
 To address the above challenges, we propose a general framework called APN.
 Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b.
-By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner.
-Based on the representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency.
+Unlike methods that resample data points within a patch, TAPA directly aggregates the raw, irregular observations.
+It learns dynamically adjustable patch boundaries to flexibly frame regions of interest.
+Crucially, it then employs a time-aware soft-weighting mechanism to generate a patch representation, ensuring that all observations are holistically considered and contribute to the representation based on their temporal relevance to the patch.
+This approach not only adapts to local information density and captures complete semantic units but also inherently prevents information loss by design, as no data point is left uncovered.
+Based on these high-quality representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency.
 Finally, predictions are made by a shallow MLP.
@@
 This framework leverages adaptive patching to generate high-quality and regular initial patch representations.
-Based on these representations, we employ a simple query module to integrate contextual information, ensuring the effective design of the framework.
+Based on these representations, we employ a simple query module to integrate contextual information, enabling a lightweight yet powerful overall architecture.
   • We design a novel TAPA module to achieve adaptive patching.
-By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequence into high-quality, regularized representations in a channel-independent manner, effectively adapting to local variations in information density and capturing complete semantics.
-  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
+By learning dynamic patch boundaries and employing a time-aware soft-aggregation scheme over raw observations, TAPA produces high-quality regularized representations without interpolation.
+This mechanism effectively adapts to local variations in information density, captures complete semantics, and guarantees full coverage of all data points.
+  • We conduct extensive experiments on multiple datasets.
+The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
+Additionally, all datasets and code are available at https://anonymous.4open.science/r/APN-23D0.
 2 Related Work
@@
 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
-Segmenting time series into “patches” to capture local patterns has demonstrated remarkable success in regular time series forecasting tasks, as evidenced by models like PatchTST (Nie et al., 2023) and Crossformer (Wang et al., 2022).
-Inspired by this, the patch-based approach has also been applied to IMTSF.
-T-PatchGNN (Zhang et al., 2024) employs its “Transformable Patch” mechanism to convert irregular sequences into patches with uniform temporal resolution but variable lengths, thereby maintaining temporal alignment while adapting to the sparsity and density variations of observation points.
-Meanwhile, TimeCHEAT (Liu et al., 2025) divides IMTS into fixed-length patches and utilizes graph learning within each patch for fine-grained temporal embedding.
-Both methods ingeniously leverage the patch strategy to transform raw IMTS data into structured representations that are more amenable to model processing, laying the groundwork for subsequent sequence modeling.
-Unlike the fixed patching methods mentioned above, this study introduces adaptive patching.
-By learning dynamically adjustable patch boundaries and employing a time-aware weighted averaging strategy, our method transforms the original irregular sequence into a high-quality regularized patch representation in a channel-independent manner.
-This approach effectively adapts to local variations in information density and captures complete semantic information.
+The patch-based strategy, which has proven successful for regular time series (Nie et al., 2023; Wang et al., 2022), has been adapted for IMTS Forecasting (IMTSF).
+Initial adaptations, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), employed a fixed-span patching strategy.
+However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS.
+It often creates patches with highly variable information content—some being too sparse for meaningful feature extraction, while others may be overly dense with redundant data.
+To overcome this, adaptive patching emerged as a natural evolution.
+One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al., 2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling.
+While effective for dense, regular data, this methodology is fundamentally flawed for IMTS.
+Interpolating across sparse or missing intervals can fabricate misleading data points, introducing significant artifacts and undermining model reliability.
+In contrast, our work proposes a distinct aggregation-based adaptive patching paradigm, specifically designed for the challenges of IMTS.
+Instead of creating new, artificial data points, our TAPA module learns dynamic "soft windows" and computes each patch representation by performing a direct, weighted aggregation of all original observations.
+This strategy provides two critical advantages for IMTS:(1)Data Fidelity: It exclusively uses the original, observed data, thereby avoiding the distortions and potential inaccuracies of interpolation.(2)Complete Information Coverage: The soft-weighting mechanism ensures that every data point contributes to the representation of relevant patches, mathematically precluding the information loss that can occur when observations fall between the hard boundaries of other methods.
 3 Methodology
@@
 3.2 Framework Overview
-The overall pipeline of APN is shown in Figure 2.
-Its core component is the Time-Aware Patch Aggregation (TAPA) Module–see Section 3.3, which processes the multivariate irregular sequence in a Channel-Independent manner.
-Compared with conventional fixed segmentation, TAPA utilizes the Adaptive Patching and Weighted Aggregation.
-This design explicitly aims to adapt to the local variations in information density inherent in IMTS and to capture complete semantic information.
-Then, a concise and efficient Query-based Aggregation module–see Section 3.4 generates a representation of historical information for each variable.
-Finally, this representation is fed into a simple MLP-based Forecasting decoder to make predictions–see Section 3.5.
-In brief, the core advantage of APN stems from the high-quality representations produced by TAPA, which effectively compresses the historical information through adapting to variations in the distribution and concentration of information and preserving complete semantic context, thus APN achieves state-of-the-art performance with a subsequent simple MLP decoder.
+The architectural design of APN is guided by a core principle: decoupling the challenge of handling irregularity from the task of forecasting.
+As illustrated in Figure 2, instead of relying on a monolithic, complex model, APN adopts a strategic two-stage pipeline.
+The cornerstone of this pipeline is the Time-Aware Patch Aggregation (TAPA) module (Section 3.3).
+Operating in a channel-independent manner, its sole yet critical purpose is to transform the raw, irregular observations into a high-quality, regularized sequence of patch representations.
+Through its novel use of Adaptive Patching and Weighted Aggregation, TAPA robustly adapts to local information density and captures complete semantic units.
+Crucially, it achieves this without resorting to interpolation, thereby preserving data fidelity and sidestepping the introduction of artificial artifacts.
+The success of this initial transformation is pivotal, as it enables the subsequent use of a remarkably efficient and lightweight architecture.
+A concise Query-based Aggregation module (Section 3.4) then effectively summarizes the historical context from these regularized patches for each variable.
+Finally, this compact representation is fed into a simple MLP-based Forecasting Decoder (Section 3.5) for the final prediction.
+In essence, APN strategically front-loads the complexity of handling irregularity into the novel TAPA module.
+By producing information-rich, regularized representations upfront, it obviates the need for computationally expensive back-end models, achieving state-of-the-art performance through an elegant and efficient design.
 3.3 Time-Aware Patch Aggregation (TAPA)
-Traditional Fixed-Span Patching (Zhang et al., 2024; Liu et al., 2025), due to the fixed patch boundaries or predefined time spans, struggles to adapt to local variations in the underlying information content and its temporal distribution (referred to as information density), and to capture complete semantic information in IMTS.
-To overcome this limitation, the TAPA mechanism devises the Adaptive Patching and Weighted Aggregation to (1) adapt the patching strategy to the local information density variations within each time series, and (2) ensure that crucial semantic information is holistically captured within the generated patches by focusing on the information most important for the prediction task.
-It efficiently integrates the information in each univariate irregular sequence $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$ by transforming it into regularized patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
-This process is performed independently for each variable, accommodating their unique sampling patterns and information distributions.
-3.3.1 Adaptive Patching
-The Adaptive Patching works through dynamically learning the temporal coverage range $[t_{p}^{left,n},t_{p}^{right,n}]$ for each target patch.
-This dynamic learning of boundaries is the primary mechanism by which TAPA adapts to local information density variations.
-For instance, in sparsely observed regions that nonetheless contain critical information, the learned patch width (controlled by $\lambda_{p}^{n}$) might expand to ensure this information is captured, or its position (controlled by $\delta_{p}^{n}$) might shift to center on these information-rich sparse points.
-Conversely, in densely observed regions with redundant information, a patch might become narrower to focus only on the novel or most salient signals, thus avoiding dilution by less informative data.
-Furthermore, this adaptive process enables the model to determine the content of each patch based on the inherent importance of the data, which is crucial for capturing complete semantic information.
-Specifically, TAPA learns the left boundary offset $\delta_{p}^{n}$ and a log-scale parameter $\lambda_{p}^{n}$ representing the patch width:
+At the heart of APN lies the Time-Aware Patch Aggregation (TAPA) module, a novel mechanism designed to fundamentally reframe how we process irregular time series.
+Traditional methods rely on a paradigm of "hard segmentation," imposing rigid, fixed-span boundaries that are ill-suited to the non-uniform nature of IMTS.
+In stark contrast, TAPA introduces a "soft aggregation" paradigm.
+Instead of discretely assigning observations to patches, it conceptualizes each patch as a soft, overlapping field of influence, allowing it to aggregate information from all observations in a temporally-aware manner.
+This paradigm shift is realized through a cohesive, two-step process executed independently for each channel $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$.
+First, in Adaptive Patching, the model learns the dynamic temporal characteristics—the center and scale—of each patch’s field of influence.
+Second, in Weighted Aggregation, it computes a representation for each patch by performing a direct, weighted aggregation of all raw observations, where the weights are determined by each observation’s temporal relevance to the patch’s learned characteristics.
+This entire process transforms the irregular input sequence into a regularized, information-rich sequence of patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
+3.3.1 Adaptive Patching: Learning Dynamic Temporal Windows
+The first step is to define the temporal scope of each patch.
+Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $[t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$.
+This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data.
+For instance, in a sparse but critical region, the model can learn to expand a window’s width to ensure sparse yet vital signals are encompassed.
+Conversely, in a dense region with redundant data, a window can be strategically narrowed to focus on the most salient signals, thus preventing dilution from less informative points.
+This adaptive process empowers the model to carve out semantically coherent segments from the raw time series, based on the data’s inherent structure.
+To achieve this, for each patch $p$ and channel $n$, TAPA learns two key parameters: a positional adjustment $\delta_{p}^{n}$ and a log-scale width parameter $\lambda_{p}^{n}$.
+The window boundaries are then computed as:
 $t_{p}^{left,n}$ $=c_{p}-\frac{S_{init}}{2}+\delta_{p}^{n},$ (2)
 $t_{p}^{right,n}$ $=t_{p}^{left,n}+\exp(\lambda_{p}^{n}),$ (3)
-where $\delta_{p}^{n}$ and $\lambda_{p}^{n}$ are learnable parameters, $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ denote the reference center points and initial average patch width.
-For observations highly important for prediction (i.e., regions with high information density relevant to the task), different patch boundaries are adjusted to accurately cover these crucial regions, ensuring their semantic integrity.
-While for regions with noise or low information relevance (low information density), boundaries may contract or shift to reduce their coverage, thus preserving the clarity of the extracted semantic information.
-The dynamic adjustment of boundaries thus constitutes an effective mechanism for adaptively selecting useful information and forming semantically coherent patches based on learned information importance.
-3.3.2 Weighted Aggregation
-To fully utilize temporal information from observations, the Learnable Time Embedding $TE(t_{i})\in\mathbb{R}^{D_{te}}$ is first generated for $i$-th timestamp, which consists of a linear layer for scale and a sine-activated linear layer for periodicity.
-Then the Learnable Time Embedding is concatenated with the original observed value $v_{i}$ to form an augmented representation $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
-After determining the dynamic coverage range $[t_{p}^{left,n},t_{p}^{right,n}]$ (which has already adapted to local information densities and aimed to cover semantic regions), we employ a query-like Weighted Aggregation.
-This weighting mechanism synergizes with the adaptive boundaries to further refine the captured semantic information by emphasizing the most salient points within each adaptively formed patch.
-For any observation $i$ (timestamp $t_{i}$) of variable $n$, its relationship $\alpha_{i,p}^{n}$ to target patch $p$ is defined by a soft window function:
-$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right),$ (4)
-where $\sigma(\cdot)$ is the Sigmoid function, and $\text{Softplus}(\kappa^{n})$ is a temperature scaling factor controlled by a learnable scalar parameter $\kappa^{n}$, regulating the smoothness of the soft window boundary transition.
-Through learning the $\kappa^{n}$, the observations within the dynamically determined patch boundaries are further weighted, where important observations contributing to high local information density receive larger $\alpha_{i,p}^{n}$ in aggregation, while less relevant observations are down-weighted or filtered out, thus ensuring the completeness and purity of the captured semantic information.
-After calculating the relationships between observations and target patches, the aggregated representation $\bar{h}_{p}^{n}$ for the $p$-th patch of variable $n$ can be obtained through a weighted average of the augmented features $\tilde{v}^{n}_{i}$:
-$\bar{h}_{p}^{n}=\frac{\sum_{i}\alpha_{i,p}^{n}\cdot\tilde{v}^{n}_{i}}{\sum_{i}\alpha_{i,p}^{n}+\epsilon}\in\mathbb{R}^{1+D_{te}},$ (5)
-where $\epsilon$ is a small positive constant to prevent division by zero.
-Subsequently, we transform $\bar{h}_{p}^{n}\in\mathbb{R}^{(1+D_{te})}$ into the uniform hidden representations $h_{p}^{n}\in\mathbb{R}^{D}$ to enhance the expressiveness through a linear projection layer $h_{p}^{n}=\text{Linear}_{D}(\bar{h}_{p}^{n})$.
-Through the aforementioned mechanisms, each univariate irregular sequence $X^{n}$ is efficiently transformed into a refined and structurally regularized patch representation sequence $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$.
+where $T_{obs}$ is the time span of the historical observation window, and $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ represent the initial reference center and width, respectively.
+The use of $\exp(\lambda_{p}^{n})$ is a deliberate design choice, ensuring the learned window width remains strictly positive throughout optimization.
+This dynamic windowing mechanism allows the model to adaptively frame regions of interest, laying the foundation for a more meaningful and robust aggregation in the subsequent step.
+3.3.2 Weighted Aggregation: From Raw Observations to Rich Representations
+With the dynamic temporal windows defined, the next step is to generate patch representations by aggregating information from the raw observations.
+Enriching Observations with Temporal Context.
+To fully leverage temporal information, we first enrich each raw observation.
+A learnable time embedding, $TE(t_{i})\in\mathbb{R}^{D_{te}}$, is generated for each timestamp $t_{i}$.
+This embedding, composed of linear layers for scale and sine-activated layers for periodicity, captures both the absolute position and periodic patterns in time.
+It is then concatenated with the original value $v_{i}$ to form an augmented representation: $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
+The Soft Window Function.
+We then employ a time-aware soft window function to calculate the relevance weight, $\alpha_{i,p}^{n}$, of each observation $i$ to each patch $p$.
+The function is elegantly constructed as the product of two Sigmoid functions:
+$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right).$ (4)
+Here, the first Sigmoid term models a smooth "falloff" from the right boundary, while the second models a smooth "rise" from the left boundary.
+Their product creates a continuous, bell-shaped weighting curve centered within the patch’s learned range.
+The term $\text{Softplus}(\kappa^{n})$ is a learnable, strictly positive temperature that controls the softness of the window edges.
+A smaller temperature leads to sharper, more defined boundaries, while a larger temperature creates gentler, more overlapping fields of influence, granting the model further adaptive flexibility.
+A Crucial Design Choice for Information Integrity.
+The choice of this soft window function is critical.
+Since the Sigmoid function $\sigma(x)$ is strictly positive for all real inputs, the resulting weight $\alpha_{i,p}^{n}$ is also guaranteed to be positive for any observation $i$ and any patch $p$.
+This property is not a mere side effect; it is a fundamental design principle that mathematically guarantees complete information coverage.
+Every observation contributes to the representation of every patch (albeit with varying degrees of influence), inherently preventing the information loss that plagues all hard-segmentation methods where observations can be inadvertently discarded.
+Final Aggregation. The final representation for the $p$-th patch, $\bar{h}_{p}^{n}$, is computed as a normalized weighted average of all augmented observation features:
+$\bar{h}_{p}^{n}=\frac{\sum_{i=1}^{L_{n}}\alpha_{i,p}^{n}\cdot\tilde{v}^{n}_{i}}{\sum_{i=1}^{L_{n}}\alpha_{i,p}^{n}+\epsilon}\in\mathbb{R}^{1+D_{te}},$ (5)
+where the summation spans all $L_{n}$ observations, and $\epsilon$ ensures numerical stability.
+Finally, to enhance expressiveness and align dimensions, we project $\bar{h}_{p}^{n}$ into the model’s uniform hidden space via a linear layer: $h_{p}^{n}=\text{Linear}_{D}(\bar{h}_{p}^{n})$.
+Through this entire process, each univariate irregular sequence is transformed into a refined, structurally regularized sequence $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$.
 3.4 Query-based Weighted Aggregation
-After TAPA generates regularized patch representation sequences $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$ for each variable, APN employs a concise mechanism to summarize them into a single representation $H_{c}^{n}$, obviating complex inter-sequence transformation modules.
-To preserve the sequential order of patches, standard positional encodings $PE\in\mathbb{R}^{P\times D}$ are added to $H^{n}$, yielding position-aware patch representations $H_{pe}^{n}=H^{n}+PE$.
-Then, to dynamically select and fuse information most relevant to the prediction task, a learnable query vector $q^{n}\in\mathbb{R}^{D}$ is introduced for each variable $n$.
-This query vector interacts with all $P$ position-aware patches $H_{pe}^{n}$, calculating unregularized importance scores $s_{p}^{n}$ via dot product, which are then regularized into weight coefficients $\beta^{n}=[\beta_{1}^{n},\dots,\beta_{P}^{n}]$ using the Softmax function.
-Finally, the representation $H_{c}^{n}$ of variable $n$ is obtained by a weighted sum of all patches $h_{pe,p}^{n}$:
+The TAPA module delivers a sequence of high-quality, regularized patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
+This successfully concludes the first stage of our pipeline: handling irregularity.
+The second stage, contextualization and forecasting, can now proceed with remarkable efficiency.
+To this end, we employ a concise query-based aggregation mechanism to distill the entire historical sequence into a single, potent context vector $H_{c}^{n}$.
+First, to preserve the sequential order of the patches, standard positional encodings ($PE$) are added, yielding position-aware representations $H_{pe}^{n}=H^{n}+PE$.
+Then, instead of resorting to complex inter-patch interactions like multi-head self-attention, we introduce a single learnable query vector $q^{n}\in\mathbb{R}^{D}$ for each channel.
+This query acts as a task-specific lens, dynamically assessing the importance of each patch for the final forecast.
+The importance scores $s_{p}^{n}$ are computed via a simple dot product, normalized into weights $\beta^{n}$ via Softmax, and used to compute the final context vector $H_{c}^{n}$ as a weighted sum:
 $s_{p}^{n}$ $=\frac{q^{n}\cdot(h_{pe,p}^{n})^{T}}{\sqrt{D}}\quad\text{for }p=1,\dots,P,$ (6)
@@
 $H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n},$ (8)
-where the summary representations of all variables $[H_{c}^{1},\dots,H_{c}^{N}]$ form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$, which is then processed by a layer normalization to stabilize subsequent MLP input.
-This learnable Query-based Weighted Aggregation is concise and efficient, avoiding complex key-value projections or pairwise inter-patch interactions, demonstrating APN’s simple yet effective design.
+The resulting summary representations for all channels, $[H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$.
+A layer normalization is then applied to stabilize the input for the decoder.
+This lightweight aggregation mechanism is a direct testament to our design philosophy: by front-loading the complexity into TAPA, the subsequent modules can be elegantly simple yet powerful.
 3.5 Forecasting Decoder
-After obtaining the summary historical representation $H_{final}^{n}\in\mathbb{R}^{D}$ for each variable $n$ and the target query time point $\tau_{k}$, APN employs a two-layer MLP decoder for prediction.
-The decoder takes the concatenation of the variable’s summary representation $H_{final}^{n}$ and the learnable temporal encoding $TE(\tau_{k})$ of the query time point $\tau_{k}$ as input, outputting the predicted value $\hat{v}^{n}_{k}$:
+The final step of the APN framework is the ultimate demonstration of its architectural elegance.
+Having distilled the entire irregular history into a powerful, fixed-size representation $H_{final}^{n}$, the forecasting task becomes remarkably straightforward.
+The prediction for a query time $\tau_{k}$ is made by a simple two-layer MLP decoder.
+This decoder takes the concatenated summary representation $H_{final}^{n}$ and the learnable temporal encoding of the query time, $TE(\tau_{k})$, as input to produce the final value $\hat{v}^{n}_{k}$:
 $\hat{v}^{n}_{k}=\text{MLP}(\text{Concat}(H_{final}^{n},TE(\tau_{k})))\in\mathbb{R}$ (9)
-3.6 Loss Function
-The overall training objective of the APN model is to minimize the Mean Squared Error (MSE) between predicted and true observed values.
-For given historical observations $\mathcal{O}_{hist}$ and a set of query time points $\mathcal{T}_{pred}=\{\tau_{k}\}_{k=1}^{H}$, the loss function $\mathcal{L}$ is defined as:
-$\mathcal{L}=\frac{1}{N\cdot H}\sum_{n=1}^{N}\sum_{k=1}^{H}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
-where $v^{n}_{k}$ is the true value of variable $n$ at query time $\tau_{k}$, and $\hat{v}^{n}_{k}$ is the model’s prediction.
+The ability to use such a simple decoder is not a limitation but a feature, underscoring the richness of the representation crafted by TAPA and the query aggregator.
+The model is trained end-to-end by minimizing the standard Mean Squared Error (MSE) loss between the predicted values $\hat{v}^{n}_{k}$ and the ground truth values $v^{n}_{k}$ across all channels and query points:
+$\mathcal{L}=\frac{1}{N\cdot Q}\sum_{n=1}^{N}\sum_{k=1}^{Q}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
+where $Q$ is the number of query points for each channel.
 4 Experiments
@@
 We train the model for a maximum of 200 epochs and employ an early stopping strategy with a patience of 50.
-All experiments are repeated 5 times, and the mean and standard deviation of the performance metrics are reported to ensure the robustness of the results.
-Table 1: Overall performance evaluated by MSE and MAE (mean $\pm$ std). The best-performing and second-best results are highlighted in bold and underline, respectively.
+To ensure the robustness of the results, we repeat each experiment 5 times using five different random seeds.
+We then report the mean and standard deviation of the performance metrics across these runs.
 Dataset PhysioNet MIMIC Human Activity USHCN
-Metric MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-2}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-1}$ MAE×$\times 10^{-1}$
+[…]
 DLinear 41.86 $\pm$ 0.05 15.52 $\pm$ 0.03 4.90 $\pm$ 0.00 16.29 $\pm$ 0.05 4.03 $\pm$ 0.01 4.21 $\pm$ 0.01 6.21 $\pm$ 0.00 3.88 $\pm$ 0.02
@@
 T-PATCHGNN 4.98 $\pm$ 0.08 3.72 $\pm$ 0.03 1.69 $\pm$ 0.03 7.22 $\pm$ 0.09 2.66 $\pm$ 0.03 3.15 $\pm$ 0.02 5.00 $\pm$ 0.04 3.08 $\pm$ 0.04
-TimeCHEAT 5.03 $\pm$ 0.07 3.88 $\pm$ 0.03 1.71 $\pm$ 0.02 7.41 $\pm$ 0.10 4.05 $\pm$ 0.10 4.64 $\pm$ 0.07 4.25 $\pm$ 0.07 3.18 $\pm$ 0.05
-APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.77 $\pm$ 0.00 3.10 $\pm$ 0.01 4.09 $\pm$ 0.06 2.92 $\pm$ 0.05
+TimeCHEAT 5.03 $\pm$ 0.07 3.88 $\pm$ 0.03 1.71 $\pm$ 0.02 7.41 $\pm$ 0.10 4.05 $\pm$ 0.10 4.64 $\pm$ 0.07 4.73 $\pm$ 0.07 3.18 $\pm$ 0.05
+APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.71 $\pm$ 0.00 3.10 $\pm$ 0.01 4.52 $\pm$ 0.06 2.92 $\pm$ 0.05
+Table 1: Overall performance evaluated by MSE and MAE (mean $\pm$ std). The best-performing and second-best results are highlighted in bold and underline, respectively.
 4.2 Main results
@@
 Whether on IMTS datasets with different characteristics in healthcare (PhysioNet, MIMIC), biomechanics (Human Activity), or climate science (USHCN), APN consistently performs excellently.
-The outstanding performance of APN can be attributed to its innovative use of adaptive patching, which generates high-quality and regular initial patch representations.
-Building on these representations, APN employs a streamlined query module to effectively integrate contextual information, ensuring a lightweight and efficient framework design.
+The outstanding performance of APN can be attributed to its TAPA module, which generates superior patch representations by directly performing a soft-aggregation over raw observations within adaptively learned temporal windows.
+This approach effectively captures salient local dynamics without the potential distortion from interpolation.
+Building on these high-quality representations, the streamlined query module and MLP decoder ensure a lightweight yet powerful framework.
 4.3 Ablation Studies
-To validate the contributions of each key component of our APN model, we conduct ablation experiments—see Table 2.
-We have the following observations: 1) w/o Adaptive Patching: Replacing the learnable boundaries in TAPA with fixed time windows leads to performance degradation across all datasets, demonstrating the importance of the adaptive boundary learning mechanism.
-2) w/o Weighted Aggregation: Replacing the time membership-based weighted averaging in TAPA with a simple arithmetic average of observations within the boundaries results in a significant performance drop.
-This highlights the core value of the time membership weighting strategy in synergizing with dynamic boundaries to further refine information and balance the information density within patches.
-3) w/o Query-Based Patch Aggregation: Replacing the query-based intra-variable patch aggregation module with a simple linear mapping also leads to a notable performance loss, indicating that even with high-quality initial patch representations, effective context information integration remains crucial.
+We perform ablation studies to isolate and validate the contribution of each primary component of APN.
+The complete results are provided in the Appendix B.2, with a representative summary presented in Table 2.
+The experiments yield three key observations.
+First, removing the Adaptive Patching mechanism and reverting to fixed-size windows results in a consistent decline in performance, which highlights the value of dynamically learned boundaries.
+Second, ablating the Weighted Aggregation scheme for a simple average over hard-cut patches leads to the most significant performance degradation.
+This empirically confirms that our soft-aggregation design is critical for preventing information loss at patch boundaries (i.e., the coverage issue).
+Third, replacing the Query-Based Patch Aggregation with a basic linear layer also diminishes performance, proving that our sophisticated aggregation method is essential for integrating contextual information within variables.
+Dataset PhysioNet MIMIC USHCN
+[…]
+w/o Adaptive
+Patching
+6.98 $\pm$ 0.05 2.05 $\pm$ 0.02 5.59 $\pm$ 0.05
+w/o Weighted
+Aggregation
+9.92 $\pm$ 0.11 2.27 $\pm$ 0.03 5.38 $\pm$ 0.02
+w/o Query-based
+Patch Aggregation
+9.92 $\pm$ 0.10 3.57 $\pm$ 0.03 5.21 $\pm$ 0.04
+APN (ours) 4.48 $\pm$ 0.06 1.61 $\pm$ 0.02 4.52 $\pm$ 0.06
 Table 2: Ablation studies for APN in terms of the best result highlighted in bold.
-Dataset PhysioNet MIMIC Human Activity USHCN
-Metric MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-2}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-1}$ MAE×$\times 10^{-1}$
-w/o Adaptive Patching 4.58 $\pm$ 0.05 3.45 $\pm$ 0.03 1.65 $\pm$ 0.02 6.94 $\pm$ 0.04 2.78 $\pm$ 0.01 3.11 $\pm$ 0.00 4.38 $\pm$ 0.05 3.07 $\pm$ 0.04
-w/o Weighted Aggregation 9.92 $\pm$ 0.11 5.81 $\pm$ 0.07 1.64 $\pm$ 0.03 6.96 $\pm$ 0.05 4.36 $\pm$ 0.12 4.28 $\pm$ 0.03 4.19 $\pm$ 0.02 2.95 $\pm$ 0.03
-w/o Query-based Patch Aggregation 9.92 $\pm$ 0.10 4.71 $\pm$ 0.09 3.57 $\pm$ 0.03 11.58 $\pm$ 0.10 2.81 $\pm$ 0.02 3.18 $\pm$ 0.05 4.51 $\pm$ 0.04 3.52 $\pm$ 0.02
-APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.77 $\pm$ 0.00 3.10 $\pm$ 0.01 4.09 $\pm$ 0.06 2.92 $\pm$ 0.05
 4.4 Parameter Sensitivity
 Figure 3: Parameter sensitivity studies of main hyper-parameters in APN
-To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Table 3.
+Figure 4: We evaluate the memory usage (GB), number of parameters (M), average training time per epoch (s), and total inference time (s) of APN and three representative baselines for IMTSF.
+All experiments are conducted on the PhysioNet dataset with a consistent batch size of 32 to ensure fair comparison.
+A lower value indicates better performance.
+To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Figure 3.
 We have the following observations: 1) From Figures 3(a) and 3(b), we observe that for most datasets, a larger $P$ does not necessarily lead to better performance.
@@
 4.5 Scalability and Efficiency Analysis
-Figure 4: We evaluate the memory usage (GB), number of parameters (M), average training time per epoch (s), and total inference time (s) of APN and three representative baselines for IMTSF.
-All experiments are conducted on the PhysioNet dataset with a consistent batch size of 32 to ensure fair comparison.
-A lower value indicates better performance.
 To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu et al., 2025), CRU (Schirmer et al., 2022), T-PatchGNN (Zhang et al., 2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4.
@@
 These representations support a highly streamlined information aggregation and prediction architecture, ultimately achieving a lightweight and highly efficient model design.
-4.6 Varying Horizon
-Table 3: Performance of varying observation and forecast horizons.
-Algorithm History=3h, Forecast=45h History=12h, Forecast=36h History=36h, Forecast=12h History=45h, Forecast=3h
-MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$
-DLinear 51.82 17.13 43.56 15.73 41.63 15.48 41.23 15.51
-TimesNet 57.30 10.70 24.95 7.62 13.57 5.50 13.86 5.65
-PatchTST 42.18 13.67 18.56 7.80 9.85 5.11 8.53 4.64
-Crossformer 9.48 5.86 8.57 5.70 5.70 4.47 5.33 4.44
-Graph Wavenet 9.43 5.86 7.23 4.82 4.71 3.90 4.10 3.73
-MTGNN 9.83 5.95 7.48 5.01 5.08 3.99 5.22 4.19
-StemGNN 8.70 5.37 7.46 4.84 6.65 4.69 5.47 4.56
-CrossGNN 10.44 6.56 7.97 5.37 6.87 4.73 4.80 4.24
-FourierGNN 9.59 5.61 7.95 4.99 6.35 4.61 5.37 4.34
-GRU-D 8.18 4.99 6.89 4.55 4.42 3.66 4.44 3.79
-SeFT 9.78 5.55 9.30 5.41 9.15 5.15 8.76 5.57
-RainDrop 10.47 5.72 9.89 5.62 9.70 5.40 9.28 5.62
-Warpformer 8.48 5.13 7.57 4.83 5.60 4.09 6.44 4.67
-mTAND 8.45 5.23 7.11 4.67 5.71 4.17 5.44 4.33
-Latent-ODE 8.25 5.04 7.20 4.69 6.70 4.36 7.10 5.33
-CRU 9.20 5.38 9.20 5.31 9.50 5.41 11.60 6.98
-Neural Flow 8.30 4.99 8.50 5.27 7.70 4.68 7.40 5.10
-T-PATCHGNN 8.01 4.87 6.48 4.19 4.14 3.31 3.69 3.25
-TimeCHEAT 9.43 5.41 6.64 4.57 5.26 4.04 2.79 3.35
-APN (ours) 8.79 4.90 5.97 3.94 3.91 3.08 2.51 2.65
-To further evaluate the robustness of APN under varying combinations of observation history lengths and forecasting horizons, we conduct experiments on the PhysioNet dataset—as shown in Table 3.
-The results demonstrate that APN exhibits strong predictive capabilities and broad adaptability across multiple "History-Forecast" settings.
-In scenarios where historical information is relatively sufficient or the prediction span is moderate, APN achieves optimal performance in both MSE and MAE metrics, leading comprehensively.
-Under extreme settings with extremely limited historical information and long prediction spans, APN’s performance is comparable to several strong baseline models, ranking second only to T-PATCHGNN.
-Considering the significant challenges faced by all models under such extreme sparsity conditions, APN’s TAPA mechanism, which relies on learning adaptive features from data, may not fully realize its potential when historical information is scarce.
-Nevertheless, APN still demonstrates competitive robustness.
 5 Conclusion
-This paper addresses the challenges of IMTS forecasting by proposing a general and efficient APN.
-The core of APN lies in its novel TAPA module.
-Through learned dynamically adjustable patch boundaries and a time-aware weighting strategy, TAPA adaptively transforms irregular sequences into a fixed number of high-quality, regularized patch representations with highly condensed information, operating in a channel-independent manner.
-Leveraging the superior initial representations generated by TAPA, APN employs an exceptionally lean query module to integrate historical information and completes predictions via a shallow MLP.
+This paper addresses the challenges of IMTS forecasting by proposing a general and efficient framework, APN.
+The core of APN lies in its novel Time-Aware Patch Aggregation (TAPA) module.
+This module introduces an aggregation-based paradigm, distinct from methods that rely on fixed patching or interpolation.
+By learning dynamic boundaries and applying a time-aware soft-weighting strategy, TAPA directly aggregates information from all raw observations to generate high-quality, regularized patch representations.
+This design choice not only ensures full data coverage and robustly handles information density variations, but also enables a streamlined and efficient overall architecture.
+Leveraging these superior representations, APN employs a concise query module and a shallow MLP to make predictions.
 Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.
-Appendix A Detailed Experimental Setup
-A.1 Dataset Details
-To comprehensively and rigorously evaluate the predictive performance of our proposed APN model, we conduct extensive experiments on four real-world Irregular Multivariate Time Series (IMTS) datasets.
-These datasets are selected for their widespread use in benchmarking and public availability, representing typical characteristics of IMTS across different domains.
-Table 4 summarizes the key statistical features of these datasets.
-For all datasets, samples are randomly partitioned into training, validation, and test sets, adhering to the standard 60%:20%:20% ratio.
-Regarding data preprocessing schemes (e.g., missing value handling, normalization methods, timestamp conversion) and prediction task configurations (e.g., historical observation window length, future prediction horizon length), we strictly follow established mainstream experimental protocols, particularly referencing benchmark settings established in related works such as T-PatchGNN [Zhang et al., 2024], to ensure the fairness and comparability of our experimental results.
-Next, we detail the characteristics of each dataset and its specific application in this study.
-Table 4: Dataset statistics.
-Dataset # Variable Avg # Obs. Avg Length Max Length
-PhysioNet 41 10.7 75 216
-MIMIC 96 1.5 46.1 643
-Human Activity 12 30.2 120.6 130
-USHCN 5 36.1 163.9 214
-  • PhysioNet [Ansari et al., 2023]: Physionet [^1] https://archive.physionet.org/challenge/2012/ dataset contains time series data collected from 12,000 ICU patients.
-Each record details 41 clinical physiological indicators during the first 48 hours after ICU admission, with irregular sampling intervals and a significant amount of missing values.
-Following the standard settings of [Ansari et al., 2023, Rubanova et al., 2019], we use the observation data from the first 24 hours after each patient’s admission to predict their physiological indicator values for the subsequent 24 hours.
-Original data and access methods are available on the official PhysioNet website.
-  • MIMIC [Naiman et al., 2024]: MIMIC [^2] https://physionet.org/content/mimiciii/1.4/ dataset is a large, publicly available database of de-identified critical care medical information.
-Following the commonly used preprocessing procedures in studies such as [Bilos et al., 2021, Naiman et al., 2024], we filter and process records to obtain 23,457 patient entries.
-Each record contains 96 irregularly sampled clinical variables from the first 48 hours after patient admission, also exhibiting significant irregularity and missingness.
-The experimental task is set to utilize data from the first 24 hours to predict variable values within the subsequent 24 hours.
-Access to this dataset requires certification and adherence to data use agreements via PhysioNet.
-  • Human Activity [Zhang et al., 2023]: Human Activity [^3] https://archive.ics.uci.edu/dataset/196/localization+data+for+person+activity dataset comprises 12 motion-related variables (e.g., 3D acceleration, 3D gyroscope) recorded by sensors worn on different body parts (wrist, chest, ankle) of test subjects, with irregular sampling times.
-To construct IMTS samples suitable for forecasting tasks, we segment the original long sequences into clips of 4000 milliseconds and use the observation data from the first 3000 milliseconds to predict sensor readings for the next 1000 milliseconds.
-Through this process, we obtain a total of 5,400 IMTS samples to evaluate the model’s capability in biomechanical signal prediction.
-  • USHCN [Ansari et al., 2023]: The USHCN dataset [^4] https://www.osti.gov/biblio/1394920 contains daily data for 5 core climate variables (daily maximum temperature, daily minimum temperature, daily precipitation, daily snow depth, daily snowfall) recorded from 1,114 weather stations across the United States between 1996 and 2000.
-This dataset is characterized by its long time span and naturally occurring spatial and temporal irregularities and data missingness due to station maintenance, equipment failure, etc.
-Following the experimental setup of [Ansari et al., 2023, Rubanova et al., 2019], we process the data for the task of using climate records from the past 24 months to predict climate conditions for the next 1 month, generating a total of 26,736 IMTS samples to test the model’s predictive performance on long-period, sparse climate data.
-A.2 Baseline Model Details
-To construct a comprehensive and convincing benchmark for the IMTS forecasting task, we integrate nineteen relevant baseline models for meticulous comparison.
-In selecting and evaluating these baseline models, we adhere to established mainstream experimental protocols [Zhang et al., 2024].
-Specifically, we adapt state-of-the-art models from regular multivariate time series (RMTS) forecasting, IMTS classification, and IMTS imputation for the current IMTS forecasting task by appropriate modifications (e.g., altering the output layer for prediction tasks and incorporating timestamps and mask information as input features).
-To ensure fairness and reliability of comparison, for some baseline models that have been rigorously evaluated in recent related works such as T-PatchGNN [Zhang et al., 2024], we prioritize referencing and reporting their publicly available optimal experimental results.
-For other models, or when evaluation under specific experimental settings is required, we reproduce them based on their official open-source implementations within a unified and consistent experimental framework.
-During reproduction, we conduct exploratory searches around the hyperparameter configurations recommended in the original papers and select the optimal parameter combinations based on performance on the validation set to report their best performance.
-This approach, combining references to existing SOTA results with careful reproduction, aims to provide a broad yet rigorous performance reference system for our APN model.
-Next, we detail the core ideas, key mechanisms, and reference code sources for each baseline model:
-Regular MTSF models:
-  • DLinear [Zeng et al., 2023]: Models the trend and periodicity/residual terms of a time series using two independent single-layer linear networks, respectively, and then sums them to obtain the final prediction.
-Its design philosophy is that simple and effective linear models remain competitive in many forecasting tasks.
-Code Reference: https://github.com/cure-lab/LTSF-Linear
-  • TimesNet [Wu et al., 2023]: Innovatively transforms one-dimensional time series into two-dimensional tensors to capture multi-periodicity.
-It identifies major periods via Fourier transform and reshapes sub-sequences of different periods into 2D tensors (with period-based "columns" and intra-period "rows"), then uses a 2D convolutional network to extract features, aiming to uniformly handle series with complex periodicities.
-Code Reference: https://github.com/thuml/TimesNet
-  • PatchTST [Nie et al., 2023]: Segments each univariate time series into fixed-length non-overlapping patches, which serve as input tokens for a Transformer.
-It employs a Channel Independence strategy, modeling each variable independently, and learns temporal dependencies through a shared Transformer backbone.
-Code Reference: https://github.com/yuqinie/PatchTST
-  • Crossformer [Wang et al., 2022]: A Transformer variant with a two-stage attention mechanism.
-The first stage captures intra-sequence temporal dependencies through dimension embedding and temporal self-attention; the second stage explicitly models interactions between different variables (dimensions) via an innovative Dimension-Segment-Wise Attention.
-Code Reference: https://github.com/Thinklab-AI/Crossformer
-  • Graph Wavenet [Wu et al., 2019]: Combines Graph Convolutional Networks (GCN) with stacked dilated causal convolutions.
-It uses an adaptive adjacency matrix to learn dynamic spatial dependencies between variables and efficiently captures long-range temporal dependencies through dilated convolutions.
-Code Reference: https://github.com/nnzhan/Graph-WaveNet
-  • MTGNN [Wu et al., 2020]: Proposes a general framework comprising a graph learning module, a graph convolution module, and a temporal convolution module.
-Its graph learning module adaptively infers unidirectional relationships between variables through node embedding and metric learning.
-Code Reference: https://github.com/nnzhan/MTGNN
-  • StemGNN [Cao et al., 2020]: First transforms spatio-temporal graph data into the spectral domain (via Graph Fourier Transform and Discrete Fourier Transform), then designs modules in the spectral domain to capture spatial and temporal dependencies, and finally obtains predictions through inverse transformation.
-Code Reference: https://github.com/microsoft/StemGNN
-  • CrossGNN [Huang et al., 2023]: To handle noise in multivariate time series, this model first generates multiple time series copies at different noise levels (data augmentation), then performs information interaction and feature refinement via cross-scale GNN and cross-variable GNN.
-Code Reference: https://github.com/hqh0728/CrossGNN
-  • FourierGNN [Yi et al., 2023]: Treats multivariate time series from a pure graph perspective, where each time series is a node on the graph, and different time points are node features.
-Graph neural operations are performed in the Fourier domain to achieve prediction.
-Code Reference: https://github.com/aikunyi/FourierGNN
-IMTS classification/imputation models:
-  • GRU-D [Che et al., 2018]: An improved GRU model that models the influence of observation intervals through an introduced learnable time decay term and handles missing values via feature-level imputation to dynamically update hidden states.
-Code Reference: https://github.com/zhiyongc/GRU-D
-  • SeFT (Set Functions for Time Series) [Horn et al., 2020]: Treats observation points of an irregular time series (typically as (time, value, variable ID) tuples) as an unordered set and models them using set functions like Deep Sets, naturally handling irregularity and variable length.
-Code Reference: https://github.com/mims-harvard/Raindrop
-  • RainDrop [Zhang et al., 2022]: Models irregularly sampled data from a sensor network as a graph, where each observation point is a node.
-It captures spatio-temporal dependencies via graph attention networks and temporal attention mechanisms, generating interpolations or predictions for arbitrary query time points.
-Code Reference: https://github.com/mims-harvard/Raindrop
-  • Warpformer [Zhang et al., 2023]: A multi-scale modeling approach for irregular clinical time series.
-Its core is the introduction of a learnable dynamic time warping module that adaptively aligns time series of different scales and sampling rates to a reference timeline before feeding them into a Transformer for representation learning.
-Code Reference: https://github.com/imJiawen/Warpformer
-  • mTAND [Shukla and Marlin, 2021]: A multi-time attention network designed for irregularly sampled time series.
-It learns embedding representations for continuous timestamps and generates fixed-length context representations for variable-length input sequences, used for interpolation or classification.
-Code Reference: https://github.com/reml-lab/mTAN
-IMTSF models:
-  • Latent ODEs [Rubanova et al., 2019]: Compresses historical observation sequences (potentially variable-length and irregular) into a latent initial state via an RNN encoder, then evolves this state in latent space using a Neural ODE (an ordinary differential equation parameterized by a neural network), and finally generates predictions through a decoder.
-Code Reference: https://github.com/YuliaRubanova/latent-ode
-  • CRU (Continuous Recurrent Units) [Schirmer et al., 2022]: Proposes a continuous-time recurrent unit whose hidden state evolution is driven by a linear stochastic differential equation, with irregularly spaced observation updates handled via Kalman filtering theory.
-Code Reference: https://github.com/boschresearch/Continuous-Recurrent-Units
-  • Neural Flows [Bilos et al., 2021]: Reformulates the solving process of Neural ODEs as a continuous normalizing flow, aiming to improve numerical solution efficiency and stability by directly modeling the solution function of the ODE rather than its derivative.
-Code Reference: https://github.com/mbilos/neural-flows-experiments
-  • T-PATCHGNN [Zhang et al., 2024]: Specifically designed for IMTS forecasting.
-Its core is to transform irregular univariate sequences into time-aligned patch sequences via a transformable patching method, then combine Transformers to capture intra-patch temporal dependencies and utilize time-adaptive graph neural networks to model dynamic inter-variable dependencies.
-Code Reference: https://github.com/usail-hkust/t-PatchGNN
-  • TimeCHEAT [Liu et al., 2025]: Proposes a channel harmony strategy, employing channel-dependent graph convolutions within patches and channel-independent Transformers between patches to balance information processing across different channels.
-Code Reference: https://github.com/Alrash/TimeCHEAT
-For all baseline models, we endeavor to use the official implementations provided by their authors or widely recognized community implementations.
-Regarding hyperparameter settings, we first refer to the recommendations in the original papers and then adjust them based on performance on the validation set to ensure fair comparison.
-Appendix B Limitation
-Although the APN method demonstrates its effectiveness, several aspects warrant further exploration:
-  • Limited Scope of Validation Datasets: APN’s current validation primarily focuses on specific types of real-world datasets.
-Its adaptability and performance on more diverse domains, such as high-frequency financial data, industrial data with strong periodicities, or traffic flow with complex covariates, require further investigation to fully assess its generalization capabilities.
-  • Performance in Extreme Scenarios: In situations with exceptionally short observation histories and very long prediction horizons, APN’s performance can be suboptimal.
-This is potentially because the data-driven adaptive mechanism struggles to converge effectively with extremely scarce information.
-Future work could explore incorporating prior knowledge or regularization to enhance robustness under such data-sparse conditions.

### v3

@@
 1 Introduction
-Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao et al., 2018; Brouwer et al., 2019; Kidger et al., 2020; Ham et al., 2019; Vio et al., 2013).
+Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao, Bi, and Chen 2018; Brouwer et al.
+2019; Kidger et al.
+2020; Ham, Kim, and Luo 2019; Vio, Díaz-Trigo, and Andreani 2013).
 Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making.
-However, the inherent irregularity of observations and missing data (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
-To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), have adopted the Fixed Patching approach—see Figure 1a.
+However, the inherent irregularity of observations and missing data (Yalavarthi et al.
+2024; Zhang et al.
+2024) in IMTS pose significant challenges to IMTSF modeling.
+To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al.
+2024) and TimeCHEAT (Liu, Cao, and Chen 2025), have adopted the Fixed Patching approach—see Figure 1a.
 This approach divides the time series into fixed-length patches with equal intervals among them.
@@
 Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times.
-For example, neural ODEs-based models (Chen et al., 2018; Rubanova et al., 2019; Gravina et al., 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin, 2021; Chen et al., 2018).
-GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024).
+For example, neural ODEs-based models (Chen et al.
+2018; Rubanova, Chen, and Duvenaud 2019; Gravina et al.
+2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin 2021; Chen et al.
+2018).
+GNNs-based models (Yalavarthi et al.
+2024; Zhang et al.
+2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al.
+2021); Transformer-based models (Chen et al.
+2023; Zhang et al.
+2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al.
+2024).
 These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS.
 While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios.
-Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024b), and CycleNet (Lin et al., 2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters.
+Notably, in regular time series forecasting task, models such as DLinear (Zeng et al.
+2023), SparseTSF (Lin et al.
+2024b), and CycleNet (Lin et al.
+2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters.
 Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
@@
 This mechanism effectively adapts to local variations in information density, captures complete semantics, and guarantees full coverage of all data points.
-  • We conduct extensive experiments on multiple datasets.
-The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
-Additionally, all datasets and code are available at https://anonymous.4open.science/r/APN-23D0.
+  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
 2 Related Work
@@
 To address these characteristics, researchers have proposed various IMTSF models.
-Some models employ approaches based on continuous-time dynamics (Chen et al., 2018; Rubanova et al., 2019; Schirmer et al., 2022; Brouwer et al., 2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points.
-For instance, Neural Flows (Bilos et al., 2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers.
-GRU-ODE-Bayes (Brouwer et al., 2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism.
-Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al., 2024; Zhang et al., 2024; Liu et al., 2025) to capture complex dependencies in IMTS.
-For example, GraFITi (Yalavarthi et al., 2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs.
-T-PatchGNN (Zhang et al., 2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
+Some models employ approaches based on continuous-time dynamics (Chen et al.
+2018; Rubanova, Chen, and Duvenaud 2019; Schirmer et al.
+2022; Brouwer et al.
+2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points.
+For instance, Neural Flows (Bilos et al.
+2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers.
+GRU-ODE-Bayes (Brouwer et al.
+2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism.
+Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al.
+2024; Zhang et al.
+2024; Liu, Cao, and Chen 2025) to capture complex dependencies in IMTS.
+For example, GraFITi (Yalavarthi et al.
+2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs.
+T-PatchGNN (Zhang et al.
+2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
-The patch-based strategy, which has proven successful for regular time series (Nie et al., 2023; Wang et al., 2022), has been adapted for IMTS Forecasting (IMTSF).
-Initial adaptations, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), employed a fixed-span patching strategy.
+The patch-based strategy, which has proven successful for regular time series (Nie et al.
+2023; Wang et al.
+2022), has been adapted for IMTS Forecasting (IMTSF).
+Initial adaptations, such as T-PatchGNN (Zhang et al.
+2024) and TimeCHEAT (Liu, Cao, and Chen 2025), employed a fixed-span patching strategy.
 However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS.
@@
 To overcome this, adaptive patching emerged as a natural evolution.
-One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al., 2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling.
+One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al.
+2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling.
 While effective for dense, regular data, this methodology is fundamentally flawed for IMTS.
@@
 This entire process transforms the irregular input sequence into a regularized, information-rich sequence of patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
-3.3.1 Adaptive Patching: Learning Dynamic Temporal Windows
+Adaptive Patching: Learning Dynamic Temporal Windows
 The first step is to define the temporal scope of each patch.
-Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $[t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$.
+Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$.
 This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data.
@@
 This dynamic windowing mechanism allows the model to adaptively frame regions of interest, laying the foundation for a more meaningful and robust aggregation in the subsequent step.
-3.3.2 Weighted Aggregation: From Raw Observations to Rich Representations
+Weighted Aggregation: From Raw Observations to Rich Representations
 With the dynamic temporal windows defined, the next step is to generate patch representations by aggregating information from the raw observations.
@@
 $H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n},$ (8)
-The resulting summary representations for all channels, $[H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$.
+The resulting summary representations for all channels, $H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$.
 A layer normalization is then applied to stabilize the input for the decoder.
@@
 Baselines: To comprehensively evaluate the performance of APN, we compare it with nineteen baseline models.
-These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), Crossformer (Wang et al., 2022), Graph Wavenet (Wu et al., 2019), MTGNN (Wu et al., 2020), StemGNN (Cao et al., 2020), CrossGNN (Huang et al., 2023), and FourierGNN (Yi et al., 2023).
-(2) IMTS classification/imputation models: including GRU-D (Che et al., 2018), SeFT (Horn et al., 2020), RainDrop (Zhang et al., 2022), Warpformer (Zhang et al., 2023), and mTAND (Shukla and Marlin, 2021).
-(3) IMTSF models: including Latent ODEs (Rubanova et al., 2019), CRU (Schirmer et al., 2022), and Neural Flows (Bilos et al., 2021), T-PATCHGNN (Zhang et al., 2024), and TimeCHEAT (Liu et al., 2025).
+These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al.
+2023), TimesNet (Wu et al.
+2023), PatchTST (Nie et al.
+2023), Crossformer (Wang et al.
+2022), Graph Wavenet (Wu et al.
+2019), MTGNN (Wu et al.
+2020), StemGNN (Cao et al.
+2020), CrossGNN (Huang et al.
+2023), and FourierGNN (Yi et al.
+2023).
+(2) IMTS classification/imputation models: including GRU-D (Che et al.
+2018), SeFT (Horn et al.
+2020), RainDrop (Zhang et al.
+2022), Warpformer (Zhang et al.
+2023), and mTAND (Shukla and Marlin 2021).
+(3) IMTSF models: including Latent ODEs (Rubanova, Chen, and Duvenaud 2019), CRU (Schirmer et al.
+2022), and Neural Flows (Bilos et al.
+2021), T-PATCHGNN (Zhang et al.
+2024), and TimeCHEAT (Liu, Cao, and Chen 2025).
 For detailed information on the baseline models, please refer to Appendix A.2.
@@
 4.5 Scalability and Efficiency Analysis
-To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu et al., 2025), CRU (Schirmer et al., 2022), T-PatchGNN (Zhang et al., 2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4.
+To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu, Cao, and Chen 2025), CRU (Schirmer et al.
+2022), T-PatchGNN (Zhang et al.
+2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4.
 The results demonstrate that APN exhibits significant advantages across all key efficiency metrics.

### v4

@@
 Abstract
-The forecasting of irregular multivariate time series (IMTS) is a critical task in domains like healthcare and climate science.
-However, this task faces two significant hurdles: 1) the inherent non-uniformity and missing data in IMTS complicate the modeling of temporal dynamics, and 2) existing methods often rely on computationally expensive architectures.
-To address these dual challenges, we introduce APN, a general and efficient forecasting framework.
-At the core of APN is a novel Time-Aware Patch Aggregation (TAPA) module that introduces an aggregation-based paradigm for adaptive patching, moving beyond the limitations of fixed-span segmentation and interpolation-based methods.
-TAPA first learns dynamic temporal boundaries to define data-driven segments.
-Crucially, instead of resampling or interpolating, it directly computes patch representations via a time-aware weighted aggregation of all raw observations, where weights are determined by each observation’s temporal relevance to the segment.
-This approach provides two key advantages: it preserves data fidelity by avoiding the introduction of artificial data points and ensures complete information coverage by design.The resulting regularized and information-rich patch representations enable the use of a lightweight query module for historical context aggregation and a simple MLP for final prediction.
-Extensive experiments on multiple real-world datasets demonstrate that APN establishes a new state-of-the-art, significantly outperforming existing methods in both prediction accuracy and computational efficiency.
+The forecasting of irregular multivariate time series (IMTS) is crucial in key areas such as healthcare, biomechanics, climate science, and astronomy.
+However, achieving accurate and practical predictions is challenging due to two main factors.
+First, the inherent irregularity and data missingness in irregular time series make modeling difficult.
+Second, most existing methods are typically complex and resource-intensive.
+In this study, we propose a general framework called APN to address these challenges.
+Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching.
+By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner.
+Additionally, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency.
+Finally, predictions are made by a shallow MLP.
+Experimental results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
 1 Introduction
-Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao, Bi, and Chen 2018; Brouwer et al.
-2019; Kidger et al.
-2020; Ham, Kim, and Luo 2019; Vio, Díaz-Trigo, and Andreani 2013).
+Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao2018; GRU-ODE; Neural-CDE; qiu2025DBLoss; qiu2025duet; wu2025k2vae; gao2025ssdts; hu2024multirc).
 Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making.
-However, the inherent irregularity of observations and missing data (Yalavarthi et al.
-2024; Zhang et al.
-2024) in IMTS pose significant challenges to IMTSF modeling.
-To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al.
-2024) and TimeCHEAT (Liu, Cao, and Chen 2025), have adopted the Fixed Patching approach—see Figure 1a.
+However, the inherent irregularity of observations and missing data (GraFITi; tPatchGNN) in IMTS poses significant challenges to IMTSF modeling.
+To address this challenge, recent IMTSF methods, such as tPatchGNN (tPatchGNN) and TimeCHEAT (TimeCHEAT), have adopted the Fixed Patching approach—see Figure 1a.
 This approach divides the time series into fixed-length patches with equal intervals among them.
@@
 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context.
-Therefore, the first challenge is to design an adaptive patching approach that can effectively adapt to local information density variations in IMTS and capture the complete semantic context.
+Therefore, the first challenge is how to design an adaptive patching approach that can adapt to the local information density variations of irregular multivariate time series and capture complete semantic information.
 Figure 1: Fixed Patching vs. Adaptive Patching.
 Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times.
-For example, neural ODEs-based models (Chen et al.
-2018; Rubanova, Chen, and Duvenaud 2019; Gravina et al.
-2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin 2021; Chen et al.
-2018).
-GNNs-based models (Yalavarthi et al.
-2024; Zhang et al.
-2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al.
-2021); Transformer-based models (Chen et al.
-2023; Zhang et al.
-2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al.
-2024).
+For example, neural-ODE-based models (Neural-ODE; Latent-ODE; TG-ODE) require computationally intensive numerical solvers to accurately model continuous-time dynamics (mTAN; Neural-ODE).
+GNN-based models (GraFITi; tPatchGNN) are hindered by the overhead of complex graph construction and multi-round node information aggregation (BadGNN); Transformer-based models (ContiFormer; Warpformer), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (BadTransformer).
 These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS.
 While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios.
-Notably, in regular time series forecasting task, models such as DLinear (Zeng et al.
-2023), SparseTSF (Lin et al.
-2024b), and CycleNet (Lin et al.
-2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters.
-Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
+Notably, in regular time series forecasting tasks, models such as SparseTSF (lin2024sparsetsf), and CycleNet (lincyclenet) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters.
+Therefore, the second challenge is how to design an efficient model for IMTSF.
 To address the above challenges, we propose a general framework called APN.
 Specifically, we design a novel Time-Aware Patch Aggregation (TAPA) module that achieves adaptive patching—see Figure 1b.
-Unlike methods that resample data points within a patch, TAPA directly aggregates the raw, irregular observations.
-It learns dynamically adjustable patch boundaries to flexibly frame regions of interest.
-Crucially, it then employs a time-aware soft-weighting mechanism to generate a patch representation, ensuring that all observations are holistically considered and contribute to the representation based on their temporal relevance to the patch.
-This approach not only adapts to local information density and captures complete semantic units but also inherently prevents information loss by design, as no data point is left uncovered.
-Based on these high-quality representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency.
+By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner.
+Based on the representations, we use a simple query module to effectively integrate historical information while maintaining the model’s efficiency.
 Finally, predictions are made by a shallow MLP.
-Results on multiple real-world datasets show that APN outperforms existing state-of-the-art methods in both efficiency and accuracy.
+Results on multiple real-world datasets show that APN outperforms existing SOTA methods in both efficiency and accuracy.
 The contributions of our paper are summarized as follows:
-  • To address the IMTSF, we propose a general framework named APN.
+  • To address IMTSF, we propose a general framework named APN.
 This framework leverages adaptive patching to generate high-quality and regular initial patch representations.
-Based on these representations, we employ a simple query module to integrate contextual information, enabling a lightweight yet powerful overall architecture.
+Based on these representations, we employ a simple query module to integrate contextual information, ensuring the effective design of the framework.
   • We design a novel TAPA module to achieve adaptive patching.
-By learning dynamic patch boundaries and employing a time-aware soft-aggregation scheme over raw observations, TAPA produces high-quality regularized representations without interpolation.
-This mechanism effectively adapts to local variations in information density, captures complete semantics, and guarantees full coverage of all data points.
-  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
+By learning dynamically adjustable patch boundaries and a time-aware weighted averaging strategy, TAPA transforms the original irregular sequences into high-quality, regularized representations in a channel-independent manner, effectively adapting to local variations in information density and capturing complete semantics.
+  • We conduct experiments on multiple datasets.
+The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
+Additionally, all datasets and code are available at: https://github.com/decisionintelligence/APN.
 2 Related Work
-2.1 Irregular Multivariate Time Series Forecasting
+2.1 Progress in Irregular Multivariate Time Series Forecasting
 IMTSF is crucial for key domains such as healthcare and climate science.
@@
 To address these characteristics, researchers have proposed various IMTSF models.
-Some models employ approaches based on continuous-time dynamics (Chen et al.
-2018; Rubanova, Chen, and Duvenaud 2019; Schirmer et al.
-2022; Brouwer et al.
-2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points.
-For instance, Neural Flows (Bilos et al.
-2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers.
-GRU-ODE-Bayes (Brouwer et al.
-2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism.
-Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al.
-2024; Zhang et al.
-2024; Liu, Cao, and Chen 2025) to capture complex dependencies in IMTS.
-For example, GraFITi (Yalavarthi et al.
-2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs.
-T-PatchGNN (Zhang et al.
-2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
+Some models employ approaches based on continuous-time dynamics (Neural-ODE; Latent-ODE; CRU; GRU-ODE), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points.
+For instance, Neural Flows (Neural-Flows) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers.
+GRU-ODE-Bayes (GRU-ODE) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism.
+Other models leverage graph neural networks and attention mechanisms (GraFITi; tPatchGNN; TimeCHEAT) to capture complex dependencies in IMTS.
+For example, GraFITi (GraFITi) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs.
+tPatchGNN (tPatchGNN) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
-The patch-based strategy, which has proven successful for regular time series (Nie et al.
-2023; Wang et al.
-2022), has been adapted for IMTS Forecasting (IMTSF).
-Initial adaptations, such as T-PatchGNN (Zhang et al.
-2024) and TimeCHEAT (Liu, Cao, and Chen 2025), employed a fixed-span patching strategy.
+The patch-based strategy, which has proven successful for regular time series (PatchTST; Crossformer; wu2025enhancing; qiu2025comprehensive; qiu2025dag; lu2025dtaf; wang2025plug), has been adapted for IMTS Forecasting (IMTSF).
+Initial adaptations, such as tPatchGNN (tPatchGNN) and TimeCHEAT (TimeCHEAT), employed a fixed-span patching strategy.
 However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS.
@@
 To overcome this, adaptive patching emerged as a natural evolution.
-One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al.
-2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling.
+One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (HDMixer)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling.
 While effective for dense, regular data, this methodology is fundamentally flawed for IMTS.
@@
 In contrast, our work proposes a distinct aggregation-based adaptive patching paradigm, specifically designed for the challenges of IMTS.
-Instead of creating new, artificial data points, our TAPA module learns dynamic "soft windows" and computes each patch representation by performing a direct, weighted aggregation of all original observations.
-This strategy provides two critical advantages for IMTS:(1)Data Fidelity: It exclusively uses the original, observed data, thereby avoiding the distortions and potential inaccuracies of interpolation.(2)Complete Information Coverage: The soft-weighting mechanism ensures that every data point contributes to the representation of relevant patches, mathematically precluding the information loss that can occur when observations fall between the hard boundaries of other methods.
+Instead of creating new, artificial data points, our TAPA module learns dynamic “soft windows” and computes each patch representation by performing a direct, weighted aggregation of all original observations.
+This strategy provides two critical advantages for IMTS: (1) Data Fidelity: It exclusively uses the original, observed data, thereby avoiding the distortions and potential inaccuracies of interpolation.
+(2) Complete Information Coverage: The soft-weighting mechanism ensures that every data point contributes to the representation of relevant patches, mathematically precluding the information loss that can occur when observations fall between the hard boundaries of other methods.
 3 Methodology
@@
 3.1 Problem Definition
-An Irregular Multivariate Time Series (IMTS) $\mathcal{O}$ consists of $N$ independent univariate sequences $\{o^{n}_{1:L_{n}}\}_{n=1}^{N}$.
+An Irregular Multivariate Time Series (IMTS) $\mathcal{O}$ consists of $N$ univariate sequences $\{o^{n}_{1:L_{n}}\}_{n=1}^{N}$.
 Each sequence $o^{n}_{1:L_{n}}$ comprises $L_{n}$ observations $(t_{i}^{n},v_{i}^{n})$, where the intervals between timestamps $t_{i}^{n}$ are irregular, and sampling across different variables is typically asynchronous.
-The IMTS forecasting task is defined as: given historical observations $\mathcal{O}$ and a set of prediction queries for all variables $\mathcal{Q}=\{[q_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$, construct and optimize a prediction model $\mathcal{F}(\cdot)$ capable of accurately predicting the future observed values $\hat{\mathcal{V}}=\{[\hat{V}_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$ corresponding to each query $q_{j}^{n}$, i.e.:
+The IMTS forecasting task is defined as: given historical observations $\mathcal{O}$ and a set of prediction queries for all variables $\mathcal{Q}=\{[q_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$, construct and optimize a prediction model $\mathcal{F}(\cdot)$ capable of accurately predicting future observed values $\hat{\mathcal{V}}=\{[\hat{v}_{j}^{n}]_{j=1}^{Q_{n}}\}_{n=1}^{N}$ corresponding to each query $q_{j}^{n}$, i.e.,
 $\mathcal{F}(\mathcal{O},\mathcal{Q})\rightarrow\hat{\mathcal{V}}$ (1)
@@
 At the heart of APN lies the Time-Aware Patch Aggregation (TAPA) module, a novel mechanism designed to fundamentally reframe how we process irregular time series.
-Traditional methods rely on a paradigm of "hard segmentation," imposing rigid, fixed-span boundaries that are ill-suited to the non-uniform nature of IMTS.
-In stark contrast, TAPA introduces a "soft aggregation" paradigm.
-Instead of discretely assigning observations to patches, it conceptualizes each patch as a soft, overlapping field of influence, allowing it to aggregate information from all observations in a temporally-aware manner.
-This paradigm shift is realized through a cohesive, two-step process executed independently for each channel $X^{n}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$.
+Traditional methods rely on a paradigm of “hard segmentation,” imposing rigid, fixed-span boundaries that are ill-suited to the non-uniform nature of IMTS.
+In stark contrast, TAPA introduces a “soft aggregation” paradigm.
+Instead of discretely assigning observations to patches, it conceptualizes each patch as a soft, overlapping field of influence, allowing it to aggregate information from all observations in a temporally aware manner.
+This paradigm shift is realized through a cohesive, two-step process executed independently for each channel $o^{n}_{1:L_{n}}=\{(t^{n}_{i},v^{n}_{i})\}_{i=1}^{L_{n}}$.
 First, in Adaptive Patching, the model learns the dynamic temporal characteristics—the center and scale—of each patch’s field of influence.
@@
 The first step is to define the temporal scope of each patch.
-Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$.
+Instead of imposing predefined, rigid boundaries, TAPA learns a dynamic temporal window, $t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$.
 This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data.
@@
 The window boundaries are then computed as:
-$t_{p}^{left,n}$ $=c_{p}-\frac{S_{init}}{2}+\delta_{p}^{n},$ (2)
+$t_{p}^{left,n}$ $=c_{p}-\frac{S_{init}}{2}+\delta_{p}^{n}$ (2)
 $t_{p}^{right,n}$ $=t_{p}^{left,n}+\exp(\lambda_{p}^{n}),$ (3)
-where $T_{obs}$ is the time span of the historical observation window, and $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ represent the initial reference center and width, respectively.
+where $T_{obs}$ is the time span of the historical observation window, with $c_{p}=(p-0.5)\cdot(T_{obs}/P)$ and $S_{init}=T_{obs}/P$ representing the initial reference center and width, respectively.
 The use of $\exp(\lambda_{p}^{n})$ is a deliberate design choice, ensuring the learned window width remains strictly positive throughout optimization.
@@
 To fully leverage temporal information, we first enrich each raw observation.
-A learnable time embedding, $TE(t_{i})\in\mathbb{R}^{D_{te}}$, is generated for each timestamp $t_{i}$.
+A learnable time embedding $TE(t_{i}^{n})\in\mathbb{R}^{D_{te}}$ is generated for each timestamp $t_{i}^{n}$ in channel $n$.
 This embedding, composed of linear layers for scale and sine-activated layers for periodicity, captures both the absolute position and periodic patterns in time.
-It is then concatenated with the original value $v_{i}$ to form an augmented representation: $\tilde{v}_{i}=\text{Concat}(v_{i},TE(t_{i}))$.
+It is then concatenated with the original value $v_{i}^{n}$ to form an augmented representation:
+$\tilde{v}_{i}^{n}=\text{Concat}(v_{i}^{n},TE(t_{i}^{n}))$ (4)
 The Soft Window Function.
@@
 The function is elegantly constructed as the product of two Sigmoid functions:
-$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right).$ (4)
-Here, the first Sigmoid term models a smooth "falloff" from the right boundary, while the second models a smooth "rise" from the left boundary.
+$\alpha_{i,p}^{n}=\sigma\left(\frac{t_{p}^{right,n}-t_{i}^{n}}{\text{Softplus}(\kappa^{n})}\right)\cdot\sigma\left(\frac{t_{i}^{n}-t_{p}^{left,n}}{\text{Softplus}(\kappa^{n})}\right)$ (5)
+Here, the first Sigmoid term models a smooth “falloff” from the right boundary, while the second models a smooth “rise” from the left boundary.
 Their product creates a continuous, bell-shaped weighting curve centered within the patch’s learned range.
@@
 This property is not a mere side effect; it is a fundamental design principle that mathematically guarantees complete information coverage.
-Every observation contributes to the representation of every patch (albeit with varying degrees of influence), inherently preventing the information loss that plagues all hard-segmentation methods where observations can be inadvertently discarded.
+Every observation contributes to the representation of every patch (albeit with varying degrees of influence), inherently preventing the information loss that plagues all hard-segmentation methods, where observations can be inadvertently discarded.
 Final Aggregation. The final representation for the $p$-th patch, $\bar{h}_{p}^{n}$, is computed as a normalized weighted average of all augmented observation features:
@@
 Through this entire process, each univariate irregular sequence is transformed into a refined, structurally regularized sequence $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$.
-3.4 Query-based Weighted Aggregation
+3.4 Query-based Aggregation
 The TAPA module delivers a sequence of high-quality, regularized patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$.
@@
 The importance scores $s_{p}^{n}$ are computed via a simple dot product, normalized into weights $\beta^{n}$ via Softmax, and used to compute the final context vector $H_{c}^{n}$ as a weighted sum:
-$s_{p}^{n}$ $=\frac{q^{n}\cdot(h_{pe,p}^{n})^{T}}{\sqrt{D}}\quad\text{for }p=1,\dots,P,$ (6)
-$\beta^{n}$ $=\text{Softmax}(\{s_{p}^{n}\}_{p=1}^{P}),$ (7)
-$H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n},$ (8)
-The resulting summary representations for all channels, $H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$.
+$s_{p}^{n}$ $=\frac{q^{n}\cdot(h_{pe,p}^{n})^{T}}{\sqrt{D}}\quad\text{for }p=1,\dots,P$ (7)
+$\beta^{n}$ $=\text{Softmax}(\{s_{p}^{n}\}_{p=1}^{P})$ (8)
+$H_{c}^{n}$ $=\sum_{p=1}^{P}\beta_{p}^{n}h_{pe,p}^{n}$ (9)
+The resulting summary representations for all channels, $H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{c}\in\mathbb{R}^{N\times D}$.
 A layer normalization is then applied to stabilize the input for the decoder.
@@
 The final step of the APN framework is the ultimate demonstration of its architectural elegance.
-Having distilled the entire irregular history into a powerful, fixed-size representation $H_{final}^{n}$, the forecasting task becomes remarkably straightforward.
-The prediction for a query time $\tau_{k}$ is made by a simple two-layer MLP decoder.
-This decoder takes the concatenated summary representation $H_{final}^{n}$ and the learnable temporal encoding of the query time, $TE(\tau_{k})$, as input to produce the final value $\hat{v}^{n}_{k}$:
-$\hat{v}^{n}_{k}=\text{MLP}(\text{Concat}(H_{final}^{n},TE(\tau_{k})))\in\mathbb{R}$ (9)
+Having distilled the entire irregular history into a powerful, fixed-size representation $H_{c}^{n}$, the forecasting task becomes remarkably straightforward.
+The prediction for a query time $q_{k}^{n}$ is made by a simple two-layer MLP decoder.
+This decoder takes the concatenated summary representation $H_{c}^{n}$ and the learnable temporal encoding of the query time, $TE(q_{k}^{n})$, as input to produce the final value $\hat{v}^{n}_{k}$:
+$\hat{v}^{n}_{k}=\text{MLP}(\text{Concat}(H_{c}^{n},TE(q_{k}^{n})))\in\mathbb{R}$ (10)
 The ability to use such a simple decoder is not a limitation but a feature, underscoring the richness of the representation crafted by TAPA and the query aggregator.
 The model is trained end-to-end by minimizing the standard Mean Squared Error (MSE) loss between the predicted values $\hat{v}^{n}_{k}$ and the ground truth values $v^{n}_{k}$ across all channels and query points:
-$\mathcal{L}=\frac{1}{N\cdot Q}\sum_{n=1}^{N}\sum_{k=1}^{Q}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
-where $Q$ is the number of query points for each channel.
+$\mathcal{L}=\frac{1}{\sum_{n=1}^{N}Q_{n}}\sum_{n=1}^{N}\sum_{k=1}^{Q_{n}}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (11)
+where $Q_{n}$ is the number of query points for each channel.
 4 Experiments
+Dataset # Variables # samples Avg # Obs. Max Length
+PhysioNet 36 11,981 308.6 47
+MIMIC 96 21,250 144.6 96
+HumanActivity 12 1,359 362.2 131
+USHCN 5 1,114 313.5 337
+Table 1: Dataset statistics.
+Dataset HumanActivity USHCN PhysioNet MIMIC
+Metric MSE MAE MSE MAE MSE MAE MSE MAE
+PrimeNet 4.2507±0.0041 1.7018±0.0011 0.4930±0.0015 0.4954±0.0018 0.7953±0.0000 0.6859±0.0001 0.9073±0.0001 0.6614±0.0001
+NeuralFlows 0.1722±0.0090 0.3150±0.0094 0.2087±0.0258 0.3157±0.0187 0.4056±0.0033 0.4466±0.0027 0.6085±0.0101 0.5306±0.0066
+CRU 0.1387±0.0073 0.2607±0.0092 0.2168±0.0162 0.3180±0.0248 0.6179±0.0045 0.5778±0.0031 0.5895±0.0092 0.5151±0.0048
+mTAN 0.0993±0.0026 0.2219±0.0047 0.5561±0.2020 0.5015±0.0968 0.3809±0.0043 0.4291±0.0035 0.9408±0.1126 0.6755±0.0459
+SeFT 1.3786±0.0024 0.9762±0.0007 0.3345±0.0022 0.4083±0.0084 0.7721±0.0021 0.6760±0.0029 0.9230±0.0015 0.6628±0.0008
+GNeuralFlow 0.3936±0.1585 0.4541±0.0841 0.2205±0.0421 0.3286±0.0412 0.8207±0.0310 0.6759±0.0100 0.8957±0.0209 0.6450±0.0072
+GRU-D 0.1893±0.0627 0.3253±0.0485 0.2097±0.0493 0.3045±0.0305 0.3419±0.0029 0.3992±0.0011 0.4759±0.0100 0.4526±0.0055
+Raindrop 0.0916±0.0072 0.2114±0.0072 0.2035±0.0336 0.3029±0.0264 0.3478±0.0019 0.4044±0.0020 0.6754±0.1829 0.5444±0.0868
+Warpformer 0.0449±0.0010 0.1228±0.0018 0.1888±0.0598 0.2939±0.0591 0.3056±0.0011 0.3661±0.0016 0.4302±0.0035 0.4025±0.0014
+tPatchGNN 0.0443±0.0009 0.1247±0.0031 0.1885±0.0403 0.3084±0.0479 0.3133±0.0053 0.3697±0.0049 0.4431±0.0115 0.4077±0.0088
+GraFITi 0.0437±0.0005 0.1221±0.0017 0.1691±0.0093 0.2777±0.0248 0.3075±0.0015 0.3637±0.0036 0.4359±0.0455 0.4142±0.0297
+APN (Ours) 0.0421±0.0001 0.1159±0.0006 0.1590±0.0137 0.2611±0.0167 0.3093±0.0011 0.3650±0.0026 0.4292±0.0027 0.4016±0.0016
+Table 2: Forecasting performance on four IMTS datasets.
+Overall performance is evaluated by MSE and MAE (mean ± std).
+The best and second-best results are highlighted in bold and with an underline, respectively.
 4.1 Setup
-Datasets: To comprehensively evaluate the model’s performance, we select four widely-used IMTS datasets, including PhysioNet, MIMIC, Activity, and USHCN.
+Datasets: To evaluate the model’s performance, we select four widely used IMTS datasets, including PhysioNet, MIMIC, HumanActivity, and USHCN.
 These datasets span multiple domains such as healthcare, biomechanics, and climate science.
-More details of the benchmark datasets are included in Appendix A.1.
-Baselines: To comprehensively evaluate the performance of APN, we compare it with nineteen baseline models.
-These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al.
-2023), TimesNet (Wu et al.
-2023), PatchTST (Nie et al.
-2023), Crossformer (Wang et al.
-2022), Graph Wavenet (Wu et al.
-2019), MTGNN (Wu et al.
-2020), StemGNN (Cao et al.
-2020), CrossGNN (Huang et al.
-2023), and FourierGNN (Yi et al.
-2023).
-(2) IMTS classification/imputation models: including GRU-D (Che et al.
-2018), SeFT (Horn et al.
-2020), RainDrop (Zhang et al.
-2022), Warpformer (Zhang et al.
-2023), and mTAND (Shukla and Marlin 2021).
-(3) IMTSF models: including Latent ODEs (Rubanova, Chen, and Duvenaud 2019), CRU (Schirmer et al.
-2022), and Neural Flows (Bilos et al.
-2021), T-PATCHGNN (Zhang et al.
-2024), and TimeCHEAT (Liu, Cao, and Chen 2025).
-For detailed information on the baseline models, please refer to Appendix A.2.
-Implementation Details: We implement APN and all related experiments using the PyTorch framework on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (@ 2.90GHz) and an NVIDIA GeForce RTX 3090 GPU.
-For our APN model, we set the hidden dimension $D$ according to different datasets.
-Specifically, the hidden dimension $D$ is set to 64 for the PhysioNet and USHCN, while it is set to 16 for the Human Activity and MIMIC.
-The time encoding dimension $D_{te}$ is uniformly set to 10.
-The number of patches $P$ is adjusted based on the characteristics of the datasets: 20 for PhysioNet, 16 for MIMIC, 75 for Human Activity, and 4 for USHCN.
-The training batch size is set to 256 for all datasets.
-The model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-2}$.
-We train the model for a maximum of 200 epochs and employ an early stopping strategy with a patience of 50.
-To ensure the robustness of the results, we repeat each experiment 5 times using five different random seeds.
-We then report the mean and standard deviation of the performance metrics across these runs.
-Dataset PhysioNet MIMIC Human Activity USHCN
-[…]
-DLinear 41.86 $\pm$ 0.05 15.52 $\pm$ 0.03 4.90 $\pm$ 0.00 16.29 $\pm$ 0.05 4.03 $\pm$ 0.01 4.21 $\pm$ 0.01 6.21 $\pm$ 0.00 3.88 $\pm$ 0.02
-TimesNet 16.48 $\pm$ 0.11 6.14 $\pm$ 0.03 5.88 $\pm$ 0.08 13.62 $\pm$ 0.07 3.12 $\pm$ 0.01 3.56 $\pm$ 0.02 5.58 $\pm$ 0.05 3.60 $\pm$ 0.04
-PatchTST 12.00 $\pm$ 0.23 6.02 $\pm$ 0.14 3.78 $\pm$ 0.03 12.43 $\pm$ 0.10 4.29 $\pm$ 0.14 4.80 $\pm$ 0.09 5.75 $\pm$ 0.01 3.57 $\pm$ 0.02
-Crossformer 6.66 $\pm$ 0.11 4.81 $\pm$ 0.11 2.65 $\pm$ 0.10 9.56 $\pm$ 0.29 4.29 $\pm$ 0.20 4.89 $\pm$ 0.17 5.25 $\pm$ 0.04 3.27 $\pm$ 0.09
-Graph Wavenet 6.04 $\pm$ 0.28 4.41 $\pm$ 0.11 2.93 $\pm$ 0.09 10.50 $\pm$ 0.15 2.89 $\pm$ 0.03 3.40 $\pm$ 0.05 5.29 $\pm$ 0.04 3.16 $\pm$ 0.09
-MTGNN 6.26 $\pm$ 0.18 4.46 $\pm$ 0.07 2.71 $\pm$ 0.23 9.55 $\pm$ 0.65 3.03 $\pm$ 0.03 3.53 $\pm$ 0.03 5.39 $\pm$ 0.05 3.34 $\pm$ 0.02
-StemGNN 6.86 $\pm$ 0.28 4.76 $\pm$ 0.19 1.73 $\pm$ 0.02 7.71 $\pm$ 0.11 8.81 $\pm$ 0.37 6.90 $\pm$ 0.02 5.75 $\pm$ 0.09 3.40 $\pm$ 0.09
-CrossGNN 7.22 $\pm$ 0.36 4.96 $\pm$ 0.12 2.95 $\pm$ 0.16 10.82 $\pm$ 0.21 3.03 $\pm$ 0.10 3.48 $\pm$ 0.08 5.66 $\pm$ 0.04 3.53 $\pm$ 0.05
-FourierGNN 6.84 $\pm$ 0.35 4.65 $\pm$ 0.12 2.55 $\pm$ 0.03 10.22 $\pm$ 0.08 2.99 $\pm$ 0.02 3.42 $\pm$ 0.02 5.82 $\pm$ 0.06 3.62 $\pm$ 0.07
-GRU-D 5.59 $\pm$ 0.09 4.08 $\pm$ 0.05 1.76 $\pm$ 0.03 7.53 $\pm$ 0.09 2.94 $\pm$ 0.05 3.53 $\pm$ 0.06 5.54 $\pm$ 0.38 3.40 $\pm$ 0.28
-SeFT 9.22 $\pm$ 0.18 5.40 $\pm$ 0.08 1.87 $\pm$ 0.01 7.84 $\pm$ 0.08 12.20 $\pm$ 0.17 8.43 $\pm$ 0.07 5.80 $\pm$ 0.19 3.70 $\pm$ 0.11
-RainDrop 9.82 $\pm$ 0.08 5.57 $\pm$ 0.06 1.99 $\pm$ 0.03 8.27 $\pm$ 0.07 14.92 $\pm$ 0.14 9.45 $\pm$ 0.05 5.78 $\pm$ 0.22 3.67 $\pm$ 0.17
-Warpformer 5.94 $\pm$ 0.35 4.21 $\pm$ 0.12 1.73 $\pm$ 0.04 7.58 $\pm$ 0.13 2.79 $\pm$ 0.04 3.39 $\pm$ 0.03 5.25 $\pm$ 0.05 3.23 $\pm$ 0.05
-mTAND 6.23 $\pm$ 0.24 4.51 $\pm$ 0.17 1.85 $\pm$ 0.06 7.73 $\pm$ 0.13 3.22 $\pm$ 0.07 3.81 $\pm$ 0.07 5.33 $\pm$ 0.05 3.26 $\pm$ 0.10
-Latent-ODE 6.05 $\pm$ 0.57 4.23 $\pm$ 0.26 1.89 $\pm$ 0.19 8.11 $\pm$ 0.52 3.34 $\pm$ 0.11 3.94 $\pm$ 0.12 5.62 $\pm$ 0.03 3.60 $\pm$ 0.12
-CRU 8.56 $\pm$ 0.26 5.16 $\pm$ 0.09 1.97 $\pm$ 0.02 7.93 $\pm$ 0.19 6.97 $\pm$ 0.78 6.30 $\pm$ 0.47 6.09 $\pm$ 0.17 3.54 $\pm$ 0.18
-Neural Flow 7.20 $\pm$ 0.07 4.67 $\pm$ 0.04 1.87 $\pm$ 0.05 8.03 $\pm$ 0.19 4.05 $\pm$ 0.13 4.46 $\pm$ 0.09 5.35 $\pm$ 0.05 3.25 $\pm$ 0.05
-T-PATCHGNN 4.98 $\pm$ 0.08 3.72 $\pm$ 0.03 1.69 $\pm$ 0.03 7.22 $\pm$ 0.09 2.66 $\pm$ 0.03 3.15 $\pm$ 0.02 5.00 $\pm$ 0.04 3.08 $\pm$ 0.04
-TimeCHEAT 5.03 $\pm$ 0.07 3.88 $\pm$ 0.03 1.71 $\pm$ 0.02 7.41 $\pm$ 0.10 4.05 $\pm$ 0.10 4.64 $\pm$ 0.07 4.73 $\pm$ 0.07 3.18 $\pm$ 0.05
-APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.71 $\pm$ 0.00 3.10 $\pm$ 0.01 4.52 $\pm$ 0.06 2.92 $\pm$ 0.05
-Table 1: Overall performance evaluated by MSE and MAE (mean $\pm$ std). The best-performing and second-best results are highlighted in bold and underline, respectively.
+Table 1 summarizes the key statistical features of these datasets.
+The PhysioNet Challenge 2012 dataset provides clinical time series from the first 48 hours of ICU stays.
+MIMIC is a large database containing de-identified health data from ICU patients.
+The HumanActivity dataset consists of biomechanics data with 3D positional variables captured from subjects performing various activities.
+For climate science, the USHCN dataset includes historical meteorological data from stations across the United States.
+All datasets are partitioned into training, validation, and test sets using a standard 80%, 10%, and 10% ratio, respectively.
+Baselines: To evaluate the performance of APN, we compare it with eleven baseline models.
+These baseline models can be broadly categorized into two groups: (1) IMTS Classification/Imputation Models: including PrimeNet (PrimeNet), SeFT (SeFT), mTAN (mTAN), GRU-D (GRU_D), Raindrop (RainDrop), and Warpformer (Warpformer).
+(2) IMTS Forecasting Models: including NeuralFlows (NeuralFlows), CRU (CRU), GNeuralFlow (GNeuralFlow), tPatchGNN (tPatchGNN), GraFITi (GraFITi).
+Implementation Details: All our experiments were conducted on a server equipped with an NVIDIA A800 GPU and implemented using the PyTorch 2.6.0+cu124 framework.
+All models are trained using the Mean Squared Error (MSE) as the loss function and optimized with the AdamW optimizer.
+We set the maximum number of training epochs to 200 and employ an early stopping strategy, where training is terminated if the model’s performance on the validation set does not improve for 10 consecutive epochs.
+To ensure a fair comparison, we primarily adopted the hyperparameter settings reported in the original papers for the baseline models.
+Building on these configurations, we conducted a further comprehensive search and fine-tuning of key hyperparameters on the validation set for all models, including our proposed APN, to ensure each achieved a competitive level of performance.
+To ensure reproducibility and mitigate the effects of randomness, each experiment is run independently with five different random seeds (from 2024 to 2028), and we report the mean and standard deviation.
+Detailed hyperparameter configurations for all models are provided in our code repository.
+We do not apply the “Drop Last” trick (qiu2024tfb; qiu2025tab) to ensure a fair comparison.
 4.2 Main results
-We compare the APN model with nineteen baseline models on four challenging datasets—see Table 1.
+We compare the APN model with eleven baseline models on four challenging datasets—see Table 2.
 We have the following observations: 1) APN achieves leading prediction accuracy across the board.
 On all datasets, APN achieves optimal or highly competitive forecasting performance.
-Compared to the second-best performing model, T-PATCHGNN, APN achieves significant reductions in MSE and MAE metrics by approximately 9.63% and 5.24%, respectively.
+Compared to the second-best performing model, GraFITi, APN achieves significant reductions in MSE and MAE metrics by approximately 2.64% and 3.61%, respectively.
 2) APN demonstrates exceptional cross-domain generalization capability and robustness.
@@
 Building on these high-quality representations, the streamlined query module and MLP decoder ensure a lightweight yet powerful framework.
+Method MIMIC USHCN
+APN (Ours) 0.4292 $\pm$ 0.0027 0.1590 $\pm$ 0.0137
+w/o Adaptive
+Patching
+0.4309 $\pm$ 0.0048 0.1717 $\pm$ 0.0283
+w/o Weighted
+Aggregation
+0.4388 $\pm$ 0.0086 0.1863 $\pm$ 0.0486
+w/o Query-based
+Aggregation
+0.4672 $\pm$ 0.0051 0.1981 $\pm$ 0.0327
+Table 3: Ablation studies for APN (MSE). Results are reported as mean $\pm$ std. Best results are highlighted in bold.
 4.3 Ablation Studies
 We perform ablation studies to isolate and validate the contribution of each primary component of APN.
-The complete results are provided in the Appendix B.2, with a representative summary presented in Table 2.
-The experiments yield three key observations.
-First, removing the Adaptive Patching mechanism and reverting to fixed-size windows results in a consistent decline in performance, which highlights the value of dynamically learned boundaries.
-Second, ablating the Weighted Aggregation scheme for a simple average over hard-cut patches leads to the most significant performance degradation.
-This empirically confirms that our soft-aggregation design is critical for preventing information loss at patch boundaries (i.e., the coverage issue).
-Third, replacing the Query-Based Patch Aggregation with a basic linear layer also diminishes performance, proving that our sophisticated aggregation method is essential for integrating contextual information within variables.
-Dataset PhysioNet MIMIC USHCN
-[…]
-w/o Adaptive
-Patching
-6.98 $\pm$ 0.05 2.05 $\pm$ 0.02 5.59 $\pm$ 0.05
-w/o Weighted
-Aggregation
-9.92 $\pm$ 0.11 2.27 $\pm$ 0.03 5.38 $\pm$ 0.02
-w/o Query-based
-Patch Aggregation
-9.92 $\pm$ 0.10 3.57 $\pm$ 0.03 5.21 $\pm$ 0.04
-APN (ours) 4.48 $\pm$ 0.06 1.61 $\pm$ 0.02 4.52 $\pm$ 0.06
-Table 2: Ablation studies for APN in terms of the best result highlighted in bold.
+The experiments, summarized in Table 3, yield three key observations.
+First, replacing the Query-Based Aggregation with a basic linear layer leads to the most significant performance degradation.
+This demonstrates that a simple summarization is insufficient and that our sophisticated query mechanism is essential for dynamically identifying and prioritizing the most salient historical patterns for the forecasting task.
+Second, ablating the Weighted Aggregation scheme and reverting to a simple average over hard-cut patches also results in a substantial decline in performance.
+This empirically confirms that our soft-aggregation design is critical for preventing information loss at patch boundaries and ensuring complete information coverage.
+Finally, removing the Adaptive Patching mechanism in favor of fixed, randomly partitioned windows consistently diminishes performance, highlighting the value of learning dynamic, data-driven temporal boundaries over imposing rigid, arbitrary ones.
+Figure 3: Parameter sensitivity studies of main hyperparameters in APN.
+Figure 4: Comparison of computational efficiency for APN and three representative baselines.
+We evaluate four key metrics: (a) peak GPU memory (GB) during a single training step, (b) total number of parameters (M), (c) average training time per step (ms), and (d) average inference time per step (ms).
+All experiments are conducted on the USHCN dataset with a consistent batch size of 32 to ensure a fair comparison.
+For all metrics, a lower value indicates better performance.
 4.4 Parameter Sensitivity
-Figure 3: Parameter sensitivity studies of main hyper-parameters in APN
-Figure 4: We evaluate the memory usage (GB), number of parameters (M), average training time per epoch (s), and total inference time (s) of APN and three representative baselines for IMTSF.
-All experiments are conducted on the PhysioNet dataset with a consistent batch size of 32 to ensure fair comparison.
-A lower value indicates better performance.
 To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Figure 3.
-We have the following observations: 1) From Figures 3(a) and 3(b), we observe that for most datasets, a larger $P$ does not necessarily lead to better performance.
-This suggests that an appropriate $P$ value can better balance the capture of local information and model complexity.
-2) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$).
-In general, a smaller hidden dimension may already be sufficient to capture effective information, while an excessively large dimension can sometimes lead to a slight performance decline.
-This indicates that the choice of $D$ should be comprehensively considered based on the scale and sparsity of the dataset.
-3) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$).
-The results indicate that a moderate $D_{te}$ value can provide the model with effective temporal information.
-An excessively high time encoding dimension may introduce too much noise, interfering with the learning of core features.
-4) The APN model exhibits a certain sensitivity to the selection of these core hyperparameters, but it generally achieves competitive results within a reasonable range.
+We have the following observations: 1) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$).
+For most datasets, performance remains relatively stable across different dimensions, indicating that an excessively large dimension does not bring significant improvement.
+This suggests that a moderately sized $D$ is sufficient to capture effective information without unnecessary complexity.
+2) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$).
+The model exhibits greater sensitivity to this parameter, as performance on datasets like PhysioNet and MIMIC degrades significantly when $D_{te}$ becomes too large.
+This indicates that a moderate $D_{te}$ is optimal for providing effective temporal information, while a higher dimension may introduce noise.
+3) From Figures 3(a) and 3(b), we observe the effect of the patch number ($P$).
+For most datasets, model performance is not highly sensitive to the number of patches, suggesting that the model can robustly capture local features across a reasonable range of $P$ values.
+4) The APN model exhibits a certain sensitivity to these core hyperparameters, especially $D_{te}$, but it generally achieves competitive results within a reasonable range.
 In practical applications, appropriate tuning should be performed based on the specific characteristics of the dataset.
 4.5 Scalability and Efficiency Analysis
-To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu, Cao, and Chen 2025), CRU (Schirmer et al.
-2022), T-PatchGNN (Zhang et al.
-2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4.
+To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (GraFITi (GraFITi), CRU (CRU), tPatchGNN (tPatchGNN)) on the USHCN dataset, with the batch size uniformly set to 32—see Figure 4.
 The results demonstrate that APN exhibits significant advantages across all key efficiency metrics.
@@
 By learning dynamic boundaries and applying a time-aware soft-weighting strategy, TAPA directly aggregates information from all raw observations to generate high-quality, regularized patch representations.
-This design choice not only ensures full data coverage and robustly handles information density variations, but also enables a streamlined and efficient overall architecture.
+This design choice not only ensures full data coverage and robustly handles information density variations but also enables a streamlined and efficient overall architecture.
 Leveraging these superior representations, APN employs a concise query module and a shallow MLP to make predictions.
 Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.
+6 Acknowledgments
+This work was partially supported by the National Natural Science Foundation of China (No.62472174) and the Fundamental Research Funds for the Central Universities.